### v2

# QuEST: Low-bit Diffusion Model Quantization via Efficient Selective Finetuning

Abstract
Diffusion models have achieved remarkable success in image generation tasks, yet their practical deployment is restrained by the high memory and time consumption. While quantization paves a way for diffusion model compression and acceleration, existing methods totally fail when the models are quantized to low-bits. In this paper, we unravel three properties in quantized diffusion models that compromise the efficacy of current methods: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules. To alleviate the intensified low-bit quantization difficulty stemming from the distribution imbalance, we propose finetuning the quantized model to better adapt to the activation distribution. Building on this idea, we identify two critical types of quantized layers: those holding vital temporal information and those sensitive to reduced bit-width, and finetune them to mitigate performance degradation with efficiency. We empirically verify that our approach modifies the activation distribution and provides meaningful temporal information, facilitating easier and more accurate quantization. Our method is evaluated over three high-resolution image generation tasks and achieves state-of-the-art performance under various bit-width settings, as well as being the first method to generate readable images on full 4-bit (i.e. W4A4) Stable Diffusion. Code is been made publicly available at URL .
Machine Learning, ICML
1 Introduction
Diffusion models (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2021) have achieved great success in the field of image generation recently. Given the high capacity and flexibility of diffusion models, there still remains two challenges that limit their efficiency (Croitoru et al., 2023). The first obstacle is the time-consuming denoising process which requires hundreds or thousands of time steps that slows down the inference speed drastically. The other one is the increasing model size driven by the need of better image fidelity and higher image resolutions. Both factors contribute to considerable latency and increased computational requirements, impeding the application of diffusion models to real-world settings where both time and computational power are restricted.
Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously (Gholami et al., 2022). It aims to compress high-bit model parameters into low-bit approximations with negligible performance degradation. For example, using 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and 8$\times$ memory reduction theoretically (Liang et al., 2021). The benefits will increase consistently when the bit-widths are further lowered. Hence, low-bit quantization of diffusion models emerges as a viable approach for enhancing their efficiency.
However, existing diffusion model quantization approaches fail under low-bit settings. These methods focus either on timestep-aware calibration data construction (Shang et al., 2023; Li et al., 2023a) or quantization noise correction (He et al., 2023b), targeting at tailoring existing quantization techniques to the properties of diffusion models that are distinct from other model types such as CNNs (Pilipović et al., 2018) and ViTs (Chitty-Venkata et al., 2023). Such approaches overlook the intrinsic diffusion model mechanisms that are related with quantization, leading to inconsistencies between method deployment and model characteristics. Therefore, we intend to identify unique features in diffusion models that are associated with quantization.
In this paper, we reveal three properties of quantized diffusion models that hinder effective quantization: The activation distributions are likely to be imbalanced, where the majority of values are close to 0 but the other values are large and appear inconsistently. Conventional quantization methods tend to either approximate either the large and sparse values, inadequately estimating the numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error. Diffusion models exhibit varying functions at distinct time steps (Choi et al., 2022), such as forming object outlines at early time steps and constructing details latter. Thus, preserving accurate temporal information is important during quantization. However, current methods struggle to represent extensive time steps with limited bit-width under low-bit quantization, leading to similar time embeddings across different time steps and results in error accumulation. Diffusion models typically have complex network architectures, incorporating convolutional layers and linear layers that construct various kinds of attention blocks (Dosovitskiy et al., 2021) and residual blocks (He et al., 2015). Whereas previous works consider each model component as equally important and apply quantization uniformly, we reveal that certain layers are particularly sensitive to perturbations from quantization, while others are more resilient.
Addressing the revealed properties, we propose a novel quantization approach for diffusion models, termed QuEST (Quantization via Efficient Selective FineTuning). Confronting property , we first theoretically justify that weight finetuning can enhance model robustness towards large perturbations under low-bit settings for reducing quantization error, whereas previous methods struggle to find appropriate quantization parameters. Then we empirically show that by finetuning the model, the activation distributions are actually modified to be more amenable to quantization. Following the idea of weight finetuning, we identify two types of layers that are responsible for performance degradation: Time embedding layers exhibiting property and attention-related layers corresponding to property . Thus, we selectively and progressively finetune these small layer subsets in conjunction with all activation quantization parameters, as illustrated in Figure 1(b). With less than 10% of the total parameters involved, QuEST not only substantially enhances low-bit quantized model performance, but is also notably time-efficient and requires only supervision from the full-precision model in a data-free manner. To preserve the essential time information, we further introduce an efficient Time-aware Activation Quantizer (shown in Figure 1(a)) that can flexibly adapt to the varying activation distributions over the numerous time steps. Our contributions are summarized as follows:
  • We uncover and validate three properties that present challenges to low-bit diffusion model quantization. We also theoretically discuss why previous methods are expected to fail and justify the sufficiency of weight finetuning under low-bit settings.
  • Addressing the identified properties, we introduce QuEST, an innovative finetuning strategy that trains the model selectively and progressively, achieving low-bit quantization capability with high time and memory efficiency. An efficient activation quantizer is also proposed for effective time information estimation. Practically, our method can be executed in a data-free and parameter-efficient manner.
  • Experiments on three high-resolution image generation tasks over four models demonstrate the superiority of our method, achieving state-of-the-art performance under various bit-width settings. We are also the first method to generate reasonable images for full 4-bit Stable Diffusion.
Figure 1: Framework of the proposed method at time step $t$. $s_{l}^{t}$ represents the quantization parameters of layer $l$ at time $t$, $Q(\textbf{w}_{l})$ represents the quantized weight of layer $l$. (a) In a quantized layer, the scaling factor $s_{l}^{t}$ is updated while the others are frozen at time $t$. (b) $e(t)$ and $x_{t}$ are the input time embedding and input image representation respectively. The model layers’ weights are optimized selectively and the quantization parameters are all updated. While different selected layers are learned progressively, the corresponding weight and scaling factors are updated simultaneously using intermediate activation alignment and task loss guidance.
2 Related Work
2.1 Diffusion Model Inference
Diffusion models (Ho et al., 2020; Rombach et al., 2021) generate samples via a learned iterative denoising process. During inference, the initial input is sampled from a Gaussian distribution: $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the final output $x_{0}$ is obtained by removing noises from $x_{T}$ subsequently using:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\tilde{\boldsymbol{\mu}}_{\theta,t}(x_{t}),\tilde{\beta_{t}}\mathbf{I}).$ (1)
In this Markov Chain, $\tilde{\boldsymbol{\mu}}_{\theta,t}$ and $\tilde{\beta_{t}}$ are calculated from the model’s output. For a typical diffusion model, the denoising stage requires hundreds to thousands of iterations. This laborious sequential process makes efficient inference and high-resolution generation extremely challenging. Latent Diffusion Models (LDMs) (Rombach et al., 2021) were developed to overcome the challenges. They encode the data into a latent space and then apply the diffusion process within this space, enabling high-resolution image generation with fewer time steps. The output images are produced by decoding the latent representations. LDM’s concept and network quantization complement each other, mutually enhancing the efficiency of diffusion models.
Practically, diffusion models typically adopt a UNet architecture (Ronneberger et al., 2015), with some modules substituted by attention blocks. In a latent diffusion model, an encoder and a decoder are incorporated for projecting images into latent representations and subsequently generating images from the latent space, respectively. Usually encoders and decoders are lightweight and computationally inexpensive, so our focus is on quantizing the UNets in LDMs, in alignment with other works (as described below).
Table 1: Comparison with Different Quantization Frameworks.
Method Data Time & Memory Low-bit
Free Efficient Compatible
PTQ
QAT
Ours
2.2 Diffusion Model Quantization
Model quantization is a dominant technique for optimizing the inference memory and speed of deep learning models by reducing the precision of the tensors used in computation. Few researches were done for quantizing diffusion models and they fall into two mainstream frameworks: Quantization-Aware Training (QAT) (Jacob et al., 2017; Li et al., 2022; Xu et al., 2023; Li et al., 2023b) and Post-Training Quantization (PTQ) (Wang et al., 2020; Nahshan et al., 2021; Li et al., 2021; Wei et al., 2023a; Liu et al., 2023). QAT methods such as Q-DM (Li et al., 2023b) consider value precision as an additional optimization constraint, training all parameters from scratch with the target task objective. While it is effective for low-bit quantization, it is even more resource-intensive than training a full-precision model. Efficient-DM (He et al., 2023a) further uses Low-rank adapters (LoRA) (Hu et al., 2021) to reduce training cost. However, it introduces extra weight parameters and still requires substantial training iterations.
Post-Training Quantization methods aim to calibrate the quantization parameters based on a small calibration set. Pioneering works such as PTQ4DM (Shang et al., 2023) and Q-Diffusion (Li et al., 2023a) focus on sampling the outputs of the full-precision model under different time steps with certain probabilities to construct the calibration set. ADP-DM (Wang et al., 2023) partition time steps into different groups and optimize each group with differentiable search algorithms. PTQ-D (He et al., 2023b) decomposes the quantization noise and correct them individually with full-precision model statistics. However, these methods generally fail under 4-bit or lower bit-width. To enable low-bit compatibility with high efficiency, our method adopts the general framework of PTQ but further introduces an efficient weight fine-tuning strategy. Table 1 shows the unique capabilities of our method, where we are able to realize QAT ability with PTQ efficiency.
3 Methodology
3.1 Preliminaries
The quantization process for a single value $x$ in a vector can be formulated as:
$\hat{x}$ $=\text{clamp}(\text{round}(\frac{x}{s})+Z;q_{min},q_{max}),$ (2)
where $\hat{x}$ is the quantized integer result, $\text{round}(\cdot)$ represents rounding algorithms such as the round-to-nearest operator (Li et al., 2021) and AdaRound (Nagel et al., 2020), $s$ is referred to as the scaling factor and $Z$ is the zero-point. clamp is the function that clamps values into the range of $[q_{min},q_{max}]$, which is determined by the bit-width. Reversely, transforming the quantized values back into the full-precision form is formulated as:
$\tilde{x}=(\hat{x}-Z)*s.$ (3)
This is denoted as the dequantization process. The quantization and dequantization processes are performed on both model weights and layer outputs (also termed as ’activation’). Equation 2 indicates that the quantization error is composed of two factors: The clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship (Li et al., 2021). While previous approaches strive for an optimal balance between the two errors, they neglect the intrinsic mechanisms in quantized diffusion models. In the next section, we investigate three properties of diffusion models that relate to quantization performance and show why typical quantization methods tend to fail.
3.2 Quantization-aware Properties of Diffusion Models
Property : The activation outputs are likely to have a majority of values close to zero, with the other values being numerically large and important. Previous works (Shang et al., 2023; Li et al., 2023a) paid attention to the varying activation distribution over different time steps, but none of them focused on the activation distribution itself. As shown in Figure 2, we find that in some layers, though the majority of values are close to 0, there exists values that are relatively large and diverse. Take the bin plot on the left for example, the original activation values (blue line) ranges from [-10, 34] but with over 99% values between [-0.6, 1.7], posing difficulties in minimizing the clipping error. This phenomenon is different from the outlier distribution observed in LLMs (Lin et al., 2023; Xiao et al., 2023), where the outliers are always positive and show up at certain positions consistently (Wei et al., 2023b). Moreover, we observe that the larger values are important for generation performance. With the large values being important and the small values appearing frequently, neither of them are negligible and need to be carefully quantized at the same time. Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped large values, generating corrupted images. This phenomenon inspires us to reshape the activation distributions via model finetuning to attain more quantization-friendly distributions.
Figure 2: Illustrations of imbalanced activation distributions. We cluster activation values into uniformly distributed bins and manipulate the frequency for better visualization. In the activation distribution of the full-precision model, the majority of values cluster near zero with sporadic large values, presenting challenges for low-bit quantization. Our method reduces the large values to enable easier quantization for improved performance.
Property : Accurate temporal information is critical for quantized diffusion model performance. As shown in Equation 1, diffusion model predictions are highly correlated with time step information. Specifically, time steps are transformed and projected into time embeddings and then added onto the image features, guiding the model to denoise correctly. Table 2 shows the performance comparison of quantizing time embeddings or preserving their full-precision values in the LDM-4 of LSUN-Bedrooms (Yu et al., 2015). Under W8A8 and W4A8, quantizing the time embeddings leads to an increase of 0.81 and 1.04 (relatively 15%) in FID, respectively. This indicates that inaccurate time embedding quantization can lead to large degradation in generation performance, yielding even larger corruptions under lower bit-widths. Due to the different focus diffusion models have on image generation at different time steps, we infer that inaccurate time embeddings can cause mismatched input image and model functionality, resulting in possible oscillations in the sequence of noise removal. Thus, preserving diffusion model performance under quantization requires accurate temporal information.
Property : Different activations have different sensitivity towards decreasing bit-width. Low-bit activation quantization is often more difficult than weight quantization and remains the key challenge for diffusion model quantization. Compared with weights than can be categorized into linear layers and convolutional layers, activation types are more diverse and complicated. We hence examine the effects on model performance when quantizing different activations. Figure 3 shows how sensitive different activations are to quantization in Stable Diffusion. We quantize three different types of activations to lower bits while maintaining the others’ bit-width (8-bit). We observe that the FeedForward layer (Feng et al., 2023) activations cause generation failure at 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits. This phenomenon indicates that a small amount of activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.
Table 2: FP & Quantized Time Embedding Comparisons.
Settings FID $\downarrow$ sFID $\downarrow$
W8A8 (Quantized Time Embedding) 7.58 22.07
W8A8 (FP Time Embedding) 6.77 22.03
W8A8 (TAQuant+$\text{SLA}_{\text{TE}}$) 5.61 21.22
W4A8 (Quantized Time Embedding) 8.59 22.74
W4A8 (FP Time Embedding) 7.55 21.69
W4A8 (TAQuant+$\text{SLA}_{\text{TE}}$) 6.95 23.17
3.3 Quantization via Efficient Selective Finetuning
In this section, we first theoretically discuss the potential reasons for the failure of previous methods and the underlying rationale for weight fine-tuning. Then we introduce a finetuning method QuEST for diffusion models that can significantly boost low-bit performance with less time and memory usage. Figure 1 illustrates the proposed framework, which is a distillation-based finetuning strategy consisting of selective weight optimization and network-wise scaling factor optimization. To further address the loss of temporal information due to quantization, we propose an efficient time-aware activation quantizer.
3.3.1 Sufficiency of Finetuning
We first review the underlying theoretical foundation underpinning previous methods, which typically employ the reconstruction-based post-training quantization approach. Denote the diffusion model’s activations at time $t$ as $\mathbf{z}_{t}=[z_{1,t},z_{2,t},...,z_{n,t}]$, the task loss as $L(z_{n,t},\overline{z}_{n,t})$, where $n$ is the number of layers and $\overline{z}_{n,t}$ is the ground-truth. $L$ can be any loss function and here we use the mean squared error (MSE). We treat quantization as a type of perturbation and formulate the influence of activation quantization using Taylor expansion, assuming model weight $\mathbf{w}$ is frozen:
$\mathbb{E}[L(z_{n,t}$ $+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]$
$\approx\Delta^{T}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (4)
where $\Delta$ is the activation perturbation, $\overline{\mathbf{g}}^{(\mathbf{z})}$ is the gradient and $\overline{\mathbf{H}}^{(z_{n,t})}$ is the Hessian matrix. According to (Li et al., 2021; Yuan et al., 2022), for a well-trained model, $\overline{\mathbf{g}}^{(z_{n,t})}=\nabla_{z_{n,t}}L$ approaches 00. Thus the above equation can be simplified to:
$\mkern-18.0mu\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta=\frac{1}{2}(\tilde{z}_{n,t}-z_{n,t})^{T}\overline{\mathbf{H}}^{(z_{n,t})}(\tilde{z}_{n,t}-z_{n,t}).\mkern-18.0mu$ (5)
This leads to a rather straightforward conclusion: Quantization error can be reduced by reconstructing the model output. However, under low-bit settings, the reasoning from Equation 3.3.1 to Equation 5 is inaccurate, where the activation perturbation $\Delta$ is too large for meaningful Taylor expansion. Thus we have the following proposition:
Proposition 3.1.
Reconstruction-based post-training quantization methods may lose their theoretical guarantee due to the large value perturbations under low-bit quantization.
To resolve the above problem, we can use the following theorem by transforming $\Delta$ into a smaller perturbation $\epsilon$:
Theorem 3.2.
Given an $n$ layer diffusion model at time $t$ with quantized activations as $\tilde{\mathbf{z}}_{t}=[\tilde{z}_{1,t},\tilde{z}_{2,t},...,\tilde{z}_{n,t}]$ and $\tilde{z}_{n,t}=z_{n,t}+\Delta$, where $z_{n,t}$ is the ground-truth and $\Delta$ is the large perturbation caused by low-bit quantization. Denote the target task MSE loss as $L(\tilde{z}_{n,t},\overline{z}_{n,t})$, the quantization error can be transformed into:
$\mathbb{E}[L(z_{n,t}+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]$
[…]
where $\mathbf{w}_{n}$ is the weight for layer $n$ and $\tilde{z}_{n,t}^{i}=\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}$, $\epsilon$ is a perturbation small enough for Taylor expansion, $K$ is a constant and $\Delta=K\epsilon$.
Theorem 3.2 indicates that: To minimize the quantization error, ideally we can finetune $\mathbf{w}_{n}$ so that for any $i$, the weights can fit to the data/activation $\tilde{z}_{n-1,t}^{i}+(i-1)\epsilon$ which the full-precision counterpart may not fit, making the model more robust to varying inputs. In other words, we are finetuning the model weights for better robustness towards large input activation perturbations, therefore enabling easier quantization. We also find that the second term in Equation 3.2 can be minimized simultaneously with the first term, so that the total quantization error can be reduced via a common goal: Activation output alignment via finetuning $\mathbf{w}_{n}$.
Figure 3: Effect of decreasing different activations’ bit-width on the model performance in Stable Diffusion. FeedForward layers experience generation failures at 6 bits, while all other linear layers barely fail at 4 bits and all convolution layers only fail at 4 bits.
3.3.2 Quantization via Selective Finetuning
Motivated by the unraveled properties of quantized diffusion models and the shortcomings in existing quantization approaches, we propose an efficient finetuning strategy for low-bit diffusion model quantization. The components of our method are detailed below.
Data-Free Partial-network-wise Training. To alleviate the need of substantial training data, we propose to adopt the general setting of Post-Training Quantization and construct the calibration set in a data-free manner. By feeding sampled Gaussian noises $x_{T}$ into the full-precision model and sample iteratively over different time steps using Equation 1, we can obtain the calibration data needed for finetuning the quantized model. In practice, the number of samples generated for each time step is 256, indicating that we only have to inference the full-precision model a few times to obtain the needed number of calibration samples.
Moreover, inspired by the fact that we can view the activation quantization parameters as additional model parameters which all contribute to the final output, we propose a partial-network-wise training strategy. Different from quantization methods using layer-wise or block-wise reconstruction (Shang et al., 2023; Li et al., 2023a) that binds quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors for each time step and only train partial weight parameters (e.g. 7% of the parameters in LDM-4). As depicted in Figure 1, the majority of the model weights remain unchanged at time step $t$, and only a small subset of parameters (including layer weights and scaling factors) related to this time step are updated using the full-precision model’s outputs as ground-truth. The choices for the weights to be finetuned are discussed in the following sections. Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, our approach update the required parameters at once. In this way, we significantly save the time and memory needed for quantization. Note that the quantization parameters for the weights are not updated once initialized, in order to maintain memory efficiency during deployment.
Table 3: Unconditional image generation results of LDM-4 models on LSUN-Bedrooms/Churches 256$\times$256. ”N/A” denotes generation failure. ”FP” denotes the full precision model.
Dataset Method
Bit-width
(W/A)
Size
[…]
LSUN- Bedrooms (LDM-4) FP 32/32 1045.6 2.95
PTQ4DM 8/8 279.1 4.75
Q-Diffusion 8/8 279.1 4.53
PTQ-D 8/8 279.1 3.75
Ours 8/8 279.1 3.07
PTQ4DM 4/8 148.4 -
Q-Diffusion 4/8 148.4 5.37
PTQ-D 4/8 148.4 5.94
Ours 4/8 148.4 3.26
PTQ4DM 4/4 148.4 N/A
Q-Diffusion 4/4 148.4 N/A
PTQ-D 4/4 148.4 N/A
Ours 4/4 148.4 5.64
LSUN- Churches (LDM-8) FP 32/32 1125.4 4.02
PTQ4DM 8/8 330.6 5.54
Q-Diffusion 8/8 330.6 6.94
PTQ-D 8/8 330.6 6.40
Ours 8/8 330.6 6.55
PTQ4DM 4/8 189.9 -
Q-Diffusion 4/8 189.9 7.80
PTQ-D 4/8 189.9 7.33
Ours 4/8 189.9 7.33
PTQ4DM 4/4 189.9 N/A
Q-Diffusion 4/4 189.9 N/A
PTQ-D 4/4 189.9 N/A
Ours 4/4 189.9 11.76
Efficient Time-aware Activation Quantization. Temporal information is a unique and essential factor in diffusion models. As highlighted in (Shang et al., 2023; Li et al., 2023a), the variation of activation distribution over different time steps is fatal for conventional PTQ methods. Also, as discussed in Section 3.2, accurate time embedding quantization is critical. To discriminate between different time steps, we design the Time-aware Activation Quantizer (TAQuant) that adopts different sets of activation quantization parameters for different time steps. While these parameters consumes negligible memory and does not effect inference speed, learning independent parameters for each time step leads to immense time cost. Inspired by the fact that neighboring time steps yield similar functions in diffusion models (Choi et al., 2022), we use the simple strategy of clustering neighboring time steps so that they share the same calibration data and quantization parameters. Then the set of activation quantization parameters are formulated as:
$\mathbf{s}=\{\mathbf{s}^{t}(x_{t}),\mathbf{s}^{2t}(x_{2t}),...,\mathbf{s}^{T}(x_{T})\},$ (7)
where the time steps used for optimization are sampled uniformly, and each chosen time step’s statistics are used to represent $T/t$ adjacent ones. Though most time steps’ quantization parameters are approximated, this strategy saves the optimization time by $t$ times without compromising much performance. Section 4.3 further provides an analysis on the trade-off between time efficiency and model performance.
Selective & Progressive Layer Alignment. As investigated in Section 3.2, we identify two crucial types of activations for image generation performance and select them for weight finetuning: Time embeddings and attention-related activations (e.g. FeedForward layers in Stable Diffusion). We progressively update these components due to their distinct, non-overlapping functionalities.
Table 4: Class-conditional image generation results of the LDM-4 model on ImageNet 256$\times$256, containing 1,000 categories.
Bit-width
(W/A)
Method
Size
[…]
32/32 FP 1529.7 11.28 7.70 364.73
8/8 Q-Diffusion 428.7 10.60 9.29 350.93
PTQ-D 428.7 10.05 9.01 359.78
Ours 428.7 10.43 6.07 365.12
4/8 Q-Diffusion 237.5 9.29 9.29 336.80
PTQ-D 237.5 8.74 7.98 344.72
Ours 237.5 8.48 6.55 346.02
4/4 Q-Diffusion 237.5 N/A N/A N/A
PTQ-D 237.5 N/A N/A N/A
Ours 237.5 5.98 7.93 202.45
Table 5: Text-to-image generation results (512$\times$512) of Stable Diffusion v1.4 using COCO2014 validation set prompts.
[…]
32/32 FP 3279.1 31.50
8/8 Q-Diffusion 949.0 31.43
Ours 949.0 31.47
4/8 Q-Diffusion 539.1 31.39
Ours 539.1 31.50
4/4 Q-Diffusion 539.1 N/A
Ours 539.1 28.85
Figure 4: Generated images compared with Q-Diffusion under two bit-width types: W8A8 and W4A4, using LDM-4 in LSUN-Bedrooms. Our method not only succeed well in generating low-bit images, but also aligns more closely with the full-precision counterpart.
During a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations. This implies that the time information operates independently from the primary network flow. Thus, we first refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}^{t}$ using:
$\mathcal{L}_{\text{TE}}^{t}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||\overline{O}(e(t);\mathbf{w}_{l})-\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}^{t})||^{2}],$ (8)
where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers, $e(t)$ is the fixed time embedding for time $t$, $\overline{O}(e(t);\mathbf{w}_{l})$ is the activation output of the full-precision model representing the ground-truth, and $\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}^{t})$ is the quantized output. This loss function indicates that the chosen weight parameters are consistently updated across different time steps to ensure robustness to diverse inputs, whereas only the subset of scaling factors relevant to each specific time step are adjusted for temporal information discrimination.
Apart from time embedding layers, attention-related layers are also vital. As depicted in Figure 3, reducing activation bit-width of these layers degrades model performance the most dramatically. Denote $\mathbb{C}_{\text{A}}$ as the set containing all these layers and given the image calibration samples $x_{t}$, we optimize $\mathbf{w}_{l}$ and $\mathbf{s}^{t}$ but exclude the quantization parameters already optimized in Equation 8:
$\mkern-18.0mu\mathcal{L}_{\text{A}}^{t}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||\overline{O}(x_{t};\mathbf{w}_{:l})-\tilde{O}(x_{t};\mathbf{w}_{:l},\mathbf{s}^{t}\setminus s_{\mathbb{C}_{\text{TE}}}^{t})||^{2}],\mkern-18.0mu$ (9)
where $\mathbf{w}_{:l}$ are the weight parameters of the first $l$ layers. This strategy resembles linear probing, where only the layer directly associated with the activation output is finetuned.
Furthermore, we aim not only for each independent layer’s quantization error reduction, but also for the final generated images’ quality improvement. Therefore, the target task loss $\sum_{t}\mathcal{L}_{d}^{t}$ is introduced, where the ground-truth is approximated by the output of the full-precision counterpart. By integrating Equation 8 and Equation 9, the final objective is formulated as follows:
$\mathcal{L}$ $=\sum_{t}((\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{d}^{t})+(\mathcal{L}_{\text{A}}^{t}+\mathcal{L}_{d}^{t}))$
$=\sum_{t}(\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{\text{A}}^{t}+2\mathbb{E}[||\overline{O}(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s}^{t})||^{2}]).$ (10)
4 Experiments
4.1 Experiment Settings
To verify the effectiveness of our proposed method, we conduct experiments on three types of generation tasks: Unconditional image generation on LSUN-Bedrooms and LSUN-Churches datasets (Yu et al., 2015), class-conditional image generation on ImageNet (Deng et al., 2009), and text-to-image generation. The model architectures we quantize include LDM and Stable Diffusion (Rombach et al., 2021), and uses ”WnAm” to represent the quantization setting: n-bit weight quantization and m-bit activation quantization. DDIM samplers (Ho et al., 2020) are adopted for LDMs and the PLMS sampler (Liu et al., 2022) is used for Stable Diffusion. We generate 256 samples per time step for constructing the calibration set, and train 20 epochs over each time step. The Adam optimizer (Kingma & Ba, 2017) is adopted and the learning rate for weight finetuning and scaling factor finetuning are set as $1e^{-5}$ and $1e^{-4}$ respectively.
[…]
4.2 Experiment Results
Unconditional Generation: We evaluate the performance of our method over LDM-4 (LSUN-Bedrooms 256$\times$256) and LDM-8 (LSUN-Churches 256$\times$256) using the DDIM sampler with 200 and 500 time steps, respectively. Results are shown in table 3. We compare with methods including PTQ4DM (Shang et al., 2023), Q-Diffusion (Li et al., 2023a) and PTQ-D (He et al., 2023b). The results are evaluated using FID, where the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet (Li et al., 2023a).
Table 6: Different component comparison and efficiency comparison on LDM-4 (LSUN-Bedrooms 256 $\times$ 256).
Method
Bit-width
(W/A)
Calibration
data size
Time cost
(GPU hours)
Memory cost
Model size
[…]
FP 32/32 - - - 1045.6 2.95
PTQ (Li et al., 2023a) 4/4 5120 23.08 10334 148.4 428.8
TAQuant 4/4 5120 11.52 9822 148.4 12.56
+ $\text{SLA}_{\text{TE}}$ 4/4 5120 13.13 12134 148.4 6.41
+ $\text{SLA}_{\text{A}}$ 4/4 5120 15.25 12450 148.4 5.78
Class-conditional Generation: We evaluate the performance using LDM-4 on ImageNet 256$\times$256 using the DDIM sampler (20 steps). As shown in Table 4, three metrics are used for evaluation. Note that sFID uses additional intermediate spatial features for calculation compared with FID. We can also see that FID is not a valid metric for ImageNet LDM-4 evaluation: All methods have lower FID when quantized to lower-bits, conflicting with human perception. We show that our method not only succeeds in W4A4 quantization, but also significantly improve the generation quality on higher bit settings. Under W8A8, our method is able to achieve 2.82 improvement over sFID and 8.13 increase in Inception Score. Under W4A8, we improve the SOTA PTQ method by 1.57 sFID and 8.77 IS.
Text-to-image Generation: We use Stable Diffusion v1.4 as the model for quantization with the PLMS sampler sampling 50 time steps. Table 5 shows the results. Images are generated based on the 10,000 prompts sampled from the COCO2014 (Lin et al., 2015) validation set, and CLIP Score is calculated based on the ViT-B/16 backbone. Given the limited works done on Stable Diffusion, we can only compare with Q-Diffusion and the full-precision baseline.
4.3 Experimental Analysis
Efficiency comparison and effects of different component. In Table 6, we present a comparison of efficiency between our method and the post-training quantization approach on the LSUN-Bedrooms dataset. Our approach, while utilizing the same quantity of calibration data as the PTQ method, is more time-efficient and incurs only a 20% increase in memory usage. We also show how each proposed component improves the generation performance. The results indicate that sequentially finetuning time embedding layers followed by attention-related layers leads to consistent improvements in model performance.
Table 7 presents the comparison of with and without the use of task loss $\mathcal{L}_{d}$. The results show that using the difference between the outputs of the quantized model and the full-precision counterpart for supervision is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for $\text{SLA}_{\text{TE}}$ and $\text{SLA}_{\text{A}}$, respectively. However, when the learning process is only supervised by the task loss, the performance degrades by 7.13 FID and 9.39 sFID for $\text{SLA}_{\text{TE}}$, indicating that the task loss alone is insufficient.
Table 7: Influence of task loss supervision on model performance.
Method
$\text{SLA}_{\text{TE}}$
w/o $\mathcal{L}_{d}$
$\text{SLA}_{\text{TE}}$
w/ $\mathcal{L}_{d}$
$\text{SLA}_{\text{A}}$
w/o $\mathcal{L}_{d}$
$\text{SLA}_{\text{A}}$
w/ $\mathcal{L}_{d}$
[…]
Time step sampling frequency analysis. While time steps are sampled uniformly, will the performance consistently improve when we sample more frequently? Table 8 illustrates the trade-off between the increased number of sampled time steps (’T-Num’) and generation performance. Metrics are calculated using 10,000 samples generated from LDM-4 on LSUN-Bedrooms. The results show that the though performance improves consistently, the amount of growth gradually decreases. Starting from 5 time steps, using 15 more time steps leads to 25.22 sFID improvement. But using another 30 time steps (from 20 to 50) only leads to 2.92 sFID improvement. Given the linearly increasing time cost associated with using more time steps, we opt for only 20 out of 200 time steps to enhance quantization efficiency.
Table 8: Effect of number of uniformly sampled time steps for W4A4 LDM-4 on LSUN-Bedrooms dataset.
T-Num 5 10 20 25 40 50
FID $\downarrow$ 19.34 15.27 12.56 12.16 11.04 10.54
sFID $\downarrow$ 54.00 37.47 28.78 27.99 26.42 25.86
How QuEST influences model properties. Our approach is based on the certain characteristics observed in quantized diffusion models, hence we aim to analyze how our fine-tuning strategy counters these identified challenges. As shown in Table 2, finetuning time embedding layers leads to even better performance than the full-precision counterpart. We infer that this might be due to both the additional parameters introduced by quantization and the additional task loss supervision. Figure 2 further implies that our method adjusts the activation distribution so that it can be more easily quantized. We see that the ranges of the activation value distributions shrink from [-10, 34] to [-4, 14] and [-11, 20] to [-4, 4]. The standard deviations also decrease from 0.171 to 0.157 and from 0.073 to 0.071. This indicates a higher concentration of values around the mean, which can effectively reduce both rounding and clipping errors.
5 Conclusion
In this work, we propose QuEST, an efficient data-free finetuning framework for low-bit diffusion model quantization. Our method is motivated by the three underlying properties found in quantized diffusion models. We also theoretically demonstrate the sufficiency of finetuning, interpreting it as a way to enhance model robustness against large activation perturbations. To alleviate the performance degradation, we propose to finetune the time embedding layers and attention-related layers under the supervision of the full-precision counterpart. A time-aware activation quantizer is also introduced to handle varying time steps. Experimental results over three high-resolution image generation tasks demonstrate the effectiveness and efficiency of QuEST, achieving low-bit compatibility with less time and memory cost.

### v3

@@
 Abstract
-Diffusion models have achieved remarkable success in image generation tasks, yet their practical deployment is restrained by the high memory and time consumption.
-While quantization paves a way for diffusion model compression and acceleration, existing methods totally fail when the models are quantized to low-bits.
-In this paper, we unravel three properties in quantized diffusion models that compromise the efficacy of current methods: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules.
+The practical deployment of diffusion models still suffers from the high memory and time overhead.
+While quantization paves a way for compression and acceleration, existing methods unfortunately fail when the models are quantized to low-bits.
+In this paper, we empirically unravel three properties in quantized diffusion models that compromise the efficacy of current methods: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules.
 To alleviate the intensified low-bit quantization difficulty stemming from the distribution imbalance, we propose finetuning the quantized model to better adapt to the activation distribution.
@@
 W4A4) Stable Diffusion.
-Code is been made publicly available at URL .
-Machine Learning, ICML
+Code is available here.
 1 Introduction
-Diffusion models (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2021) have achieved great success in the field of image generation recently.
-Given the high capacity and flexibility of diffusion models, there still remains two challenges that limit their efficiency (Croitoru et al., 2023).
-The first obstacle is the time-consuming denoising process which requires hundreds or thousands of time steps that slows down the inference speed drastically.
+Diffusion models [6, 15, 32] have achieved great success in the field of image generation recently.
+Given the high capacity and flexibility of diffusion models, there still remains two challenges that limit their efficiency [4].
+The first obstacle is the time-consuming denoising process which requires hundreds to thousands of inference time steps that slows down the generation speed drastically.
 The other one is the increasing model size driven by the need of better image fidelity and higher image resolutions.
-Both factors contribute to considerable latency and increased computational requirements, impeding the application of diffusion models to real-world settings where both time and computational power are restricted.
-Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously (Gholami et al., 2022).
+Both factors contribute to considerable latency and increased computational requirements, impeding the application of diffusion models to real-world settings where both time and computational power are carefully restricted.
+Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously [9].
 It aims to compress high-bit model parameters into low-bit approximations with negligible performance degradation.
-For example, using 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and 8$\times$ memory reduction theoretically (Liang et al., 2021).
+For example, 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and 8$\times$ memory reduction theoretically [23].
 The benefits will increase consistently when the bit-widths are further lowered.
-Hence, low-bit quantization of diffusion models emerges as a viable approach for enhancing their efficiency.
+Hence, low-bit quantization of diffusion models emerges as a viable approach for efficiency enhancement.
 However, existing diffusion model quantization approaches fail under low-bit settings.
-These methods focus either on timestep-aware calibration data construction (Shang et al., 2023; Li et al., 2023a) or quantization noise correction (He et al., 2023b), targeting at tailoring existing quantization techniques to the properties of diffusion models that are distinct from other model types such as CNNs (Pilipović et al., 2018) and ViTs (Chitty-Venkata et al., 2023).
+These methods focus either on timestep-aware calibration data construction [19, 34] or quantization noise correction [12], targeting at tailoring existing quantization techniques to the properties of diffusion models that are distinct from other model types such as CNNs [31] and ViTs [2].
 Such approaches overlook the intrinsic diffusion model mechanisms that are related with quantization, leading to inconsistencies between method deployment and model characteristics.
 Therefore, we intend to identify unique features in diffusion models that are associated with quantization.
-In this paper, we reveal three properties of quantized diffusion models that hinder effective quantization: The activation distributions are likely to be imbalanced, where the majority of values are close to 0 but the other values are large and appear inconsistently.
-Conventional quantization methods tend to either approximate either the large and sparse values, inadequately estimating the numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error.
-Diffusion models exhibit varying functions at distinct time steps (Choi et al., 2022), such as forming object outlines at early time steps and constructing details latter.
+In this paper, we reveal three properties of quantized diffusion models that hinder effective quantization: The activation distributions are likely to be imbalanced, where the majority of values are close to 0 but the important and large values are rare and appear inconsistently.
+Conventional quantization methods tend to either approximate the large and sparse values, inadequately estimating the numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error.
+Diffusion models exhibit varying functions at distinct time steps [3], such as forming object outlines at early time steps and constructing details latter.
 Thus, preserving accurate temporal information is important during quantization.
-However, current methods struggle to represent extensive time steps with limited bit-width under low-bit quantization, leading to similar time embeddings across different time steps and results in error accumulation.
-Diffusion models typically have complex network architectures, incorporating convolutional layers and linear layers that construct various kinds of attention blocks (Dosovitskiy et al., 2021) and residual blocks (He et al., 2015).
+However, current methods struggle to represent extensive time steps with limited bit-width, leading to similar time embeddings across different time steps and results in error accumulation.
+Diffusion models typically have complex network architectures, incorporating convolutional layers and linear layers that construct various kinds of attention blocks [7] and residual blocks [10].
 Whereas previous works consider each model component as equally important and apply quantization uniformly, we reveal that certain layers are particularly sensitive to perturbations from quantization, while others are more resilient.
+Table 1: Comparison with different frameworks.
+Method Data Time & Memory Low-bit Fully Free Efficient Compatible Quantized PTQ [19] QAT [20] EfficientDM [11] Ours
 Addressing the revealed properties, we propose a novel quantization approach for diffusion models, termed QuEST (Quantization via Efficient Selective FineTuning).
-Confronting property , we first theoretically justify that weight finetuning can enhance model robustness towards large perturbations under low-bit settings for reducing quantization error, whereas previous methods struggle to find appropriate quantization parameters.
-Then we empirically show that by finetuning the model, the activation distributions are actually modified to be more amenable to quantization.
-Following the idea of weight finetuning, we identify two types of layers that are responsible for performance degradation: Time embedding layers exhibiting property and attention-related layers corresponding to property .
-Thus, we selectively and progressively finetune these small layer subsets in conjunction with all activation quantization parameters, as illustrated in Figure 1(b).
+Confronting property , we first theoretically justify that weight finetuning can enhance model robustness towards large perturbations in activations under low-bit settings for reducing quantization error, whereas previous methods struggle to find appropriate quantization parameters.
+Then we empirically show that by finetuning the model, the activation distributions are actually modified to be more amenable to quantization by reducing the significant values (large yet sparse), as shown in Figure 1(a).
+Following the idea of weight finetuning, we compare different bit-width models (Figure 1(b)) and identify two types of layers that are responsible for performance degradation: Time embedding layers exhibiting property and attention-related layers corresponding to property .
+Thus, we selectively and progressively finetune these small layer subsets in conjunction with all activation quantization parameters, as illustrated in Figure 1(c).
 With less than 10% of the total parameters involved, QuEST not only substantially enhances low-bit quantized model performance, but is also notably time-efficient and requires only supervision from the full-precision model in a data-free manner.
-To preserve the essential time information, we further introduce an efficient Time-aware Activation Quantizer (shown in Figure 1(a)) that can flexibly adapt to the varying activation distributions over the numerous time steps.
 Our contributions are summarized as follows:
-  • We uncover and validate three properties that present challenges to low-bit diffusion model quantization.
+  • We uncover and validate three properties that present challenges to low-bit diffusion model quantization: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules.
 We also theoretically discuss why previous methods are expected to fail and justify the sufficiency of weight finetuning under low-bit settings.
-  • Addressing the identified properties, we introduce QuEST, an innovative finetuning strategy that trains the model selectively and progressively, achieving low-bit quantization capability with high time and memory efficiency.
-An efficient activation quantizer is also proposed for effective time information estimation.
-Practically, our method can be executed in a data-free and parameter-efficient manner.
+  • Addressing the identified properties, we introduce QuEST, a new finetuning strategy that trains the model selectively and progressively, achieving low-bit quantization capability with high time and memory efficiency.
+An efficient activation quantizer is also proposed for time information estimation.
+Our method can be executed in a data-free and parameter-efficient manner.
   • Experiments on three high-resolution image generation tasks over four models demonstrate the superiority of our method, achieving state-of-the-art performance under various bit-width settings.
 We are also the first method to generate reasonable images for full 4-bit Stable Diffusion.
-Figure 1: Framework of the proposed method at time step $t$.
-$s_{l}^{t}$ represents the quantization parameters of layer $l$ at time $t$, $Q(\textbf{w}_{l})$ represents the quantized weight of layer $l$.
-(a) In a quantized layer, the scaling factor $s_{l}^{t}$ is updated while the others are frozen at time $t$.
-(b) $e(t)$ and $x_{t}$ are the input time embedding and input image representation respectively.
-The model layers’ weights are optimized selectively and the quantization parameters are all updated.
-While different selected layers are learned progressively, the corresponding weight and scaling factors are updated simultaneously using intermediate activation alignment and task loss guidance.
-2 Related Work
-2.1 Diffusion Model Inference
-Diffusion models (Ho et al., 2020; Rombach et al., 2021) generate samples via a learned iterative denoising process.
-During inference, the initial input is sampled from a Gaussian distribution: $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the final output $x_{0}$ is obtained by removing noises from $x_{T}$ subsequently using:
-$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\tilde{\boldsymbol{\mu}}_{\theta,t}(x_{t}),\tilde{\beta_{t}}\mathbf{I}).$ (1)
-In this Markov Chain, $\tilde{\boldsymbol{\mu}}_{\theta,t}$ and $\tilde{\beta_{t}}$ are calculated from the model’s output.
-For a typical diffusion model, the denoising stage requires hundreds to thousands of iterations.
-This laborious sequential process makes efficient inference and high-resolution generation extremely challenging.
-Latent Diffusion Models (LDMs) (Rombach et al., 2021) were developed to overcome the challenges.
-They encode the data into a latent space and then apply the diffusion process within this space, enabling high-resolution image generation with fewer time steps.
-The output images are produced by decoding the latent representations.
-LDM’s concept and network quantization complement each other, mutually enhancing the efficiency of diffusion models.
-Practically, diffusion models typically adopt a UNet architecture (Ronneberger et al., 2015), with some modules substituted by attention blocks.
-In a latent diffusion model, an encoder and a decoder are incorporated for projecting images into latent representations and subsequently generating images from the latent space, respectively.
-Usually encoders and decoders are lightweight and computationally inexpensive, so our focus is on quantizing the UNets in LDMs, in alignment with other works (as described below).
-Table 1: Comparison with Different Quantization Frameworks.
-Method Data Time & Memory Low-bit
-Free Efficient Compatible
-PTQ
-QAT
-Ours
-2.2 Diffusion Model Quantization
-Model quantization is a dominant technique for optimizing the inference memory and speed of deep learning models by reducing the precision of the tensors used in computation.
-Few researches were done for quantizing diffusion models and they fall into two mainstream frameworks: Quantization-Aware Training (QAT) (Jacob et al., 2017; Li et al., 2022; Xu et al., 2023; Li et al., 2023b) and Post-Training Quantization (PTQ) (Wang et al., 2020; Nahshan et al., 2021; Li et al., 2021; Wei et al., 2023a; Liu et al., 2023).
-QAT methods such as Q-DM (Li et al., 2023b) consider value precision as an additional optimization constraint, training all parameters from scratch with the target task objective.
-While it is effective for low-bit quantization, it is even more resource-intensive than training a full-precision model.
-Efficient-DM (He et al., 2023a) further uses Low-rank adapters (LoRA) (Hu et al., 2021) to reduce training cost.
-However, it introduces extra weight parameters and still requires substantial training iterations.
-Post-Training Quantization methods aim to calibrate the quantization parameters based on a small calibration set.
-Pioneering works such as PTQ4DM (Shang et al., 2023) and Q-Diffusion (Li et al., 2023a) focus on sampling the outputs of the full-precision model under different time steps with certain probabilities to construct the calibration set.
-ADP-DM (Wang et al., 2023) partition time steps into different groups and optimize each group with differentiable search algorithms.
-PTQ-D (He et al., 2023b) decomposes the quantization noise and correct them individually with full-precision model statistics.
-However, these methods generally fail under 4-bit or lower bit-width.
-To enable low-bit compatibility with high efficiency, our method adopts the general framework of PTQ but further introduces an efficient weight fine-tuning strategy.
-Table 1 shows the unique capabilities of our method, where we are able to realize QAT ability with PTQ efficiency.
+The detailed related works on diffusion model inference and quantization are referred to appendix for space restriction.
+Table 1 is a compact summary that shows the unique capabilities of our method, where we are able to overcome the limitations of previous quantization methods.
+In comparison with EfficientDM [11], we include a brief discussion here: Technically, EfficientDM aims to finetune via low-rank weight representations [16].
+However, it introduces additional parameters for each layer and requires extensive iterations of training on LoRA weights.
+Our method instead does not include additional parameters and only needs to finetune the model for limited iterations with the same batch size.
+We have also found that EfficientDM does not quantize the matrix multiplications in the attention mechanism, while we quantize the matrices to 8-bits.
+Moreover, EfficientDM always keeps the first three layers and the last layer as full 8-bit (weight and activation), as well as keeping skip-connections and some linear layers as full-precision during implementation.
+Our method instead only kept the first and last layer as 8-bit and quantize all other layers to the required bit-width.
+Figure 1: Overview of our findings and method.
+(a) Illustration of property and a way to mitigate the quantization difficulty.
+(b) Demonstration of how property and property are identified via comparing different bit-width models.
+(c) Framework of the proposed method at time step $t$.
+To address the quantization difficulty and achieve more quantization-friendly activation distribution efficiently, we reduce significance by selectively and progressively updating partial model layers’ weights.
+$e(t)$ and $x_{t}$ are the input time embedding and input image representation respectively.
+While $\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$ are selected layer weights to be optimized using Equation 2.3.2, $\mathbf{W}_{l^{{}^{\prime}}}$ and $\mathbf{W}_{l^{{}^{\prime\prime}}}$ are frozen during finetuning.
+More details are included in Section 2.3.2.
 2 Methodology
@@
 $\hat{x}$ $=\text{clamp}(\text{round}(\frac{x}{s})+Z;q_{min},q_{max}),$ (1)
-where $\hat{x}$ is the quantized integer result, $\text{round}(\cdot)$ represents rounding algorithms such as the round-to-nearest operator (Li et al., 2021) and AdaRound (Nagel et al., 2020), $s$ is referred to as the scaling factor and $Z$ is the zero-point.
+where $\hat{x}$ is the quantized integer result, $\text{round}(\cdot)$ represents rounding algorithms such as the round-to-nearest operator [22] and AdaRound [27], $s$ is referred to as the scaling factor and $Z$ is the zero-point.
 clamp is the function that clamps values into the range of $[q_{min},q_{max}]$, which is determined by the bit-width.
-Reversely, transforming the quantized values back into the full-precision form is formulated as:
+Reversely, transforming the quantized values back into the full-precision form is:
 $\tilde{x}=(\hat{x}-Z)*s.$ (2)
@@
 The quantization and dequantization processes are performed on both model weights and layer outputs (also termed as ’activation’).
-Equation 2 indicates that the quantization error is composed of two factors: The clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship (Li et al., 2021).
+Equation 1 indicates that the quantization error is composed of two factors: The clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship [22].
 While previous approaches strive for an optimal balance between the two errors, they neglect the intrinsic mechanisms in quantized diffusion models.
@@
 Property : The activation outputs are likely to have a majority of values close to zero, with the other values being numerically large and important.
-Previous works (Shang et al., 2023; Li et al., 2023a) paid attention to the varying activation distribution over different time steps, but none of them focused on the activation distribution itself.
-As shown in Figure 2, we find that in some layers, though the majority of values are close to 0, there exists values that are relatively large and diverse.
-Take the bin plot on the left for example, the original activation values (blue line) ranges from [-10, 34] but with over 99% values between [-0.6, 1.7], posing difficulties in minimizing the clipping error.
-This phenomenon is different from the outlier distribution observed in LLMs (Lin et al., 2023; Xiao et al., 2023), where the outliers are always positive and show up at certain positions consistently (Wei et al., 2023b).
-Moreover, we observe that the larger values are important for generation performance.
-With the large values being important and the small values appearing frequently, neither of them are negligible and need to be carefully quantized at the same time.
-Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped large values, generating corrupted images.
-This phenomenon inspires us to reshape the activation distributions via model finetuning to attain more quantization-friendly distributions.
-Figure 2: Illustrations of imbalanced activation distributions.
-We cluster activation values into uniformly distributed bins and manipulate the frequency for better visualization.
+Previous works [19, 34] paid attention to the varying activation distribution over different time steps, but none of them focused on the activation distribution itself.
+In Figure 2, we cluster activation values into uniformly distributed bins and manipulate the frequency for better visualization.
+We find that in some layers, though the majority of values are close to 0, there exists values that are relatively large and diverse, which we term as "significant values" (circled in green).
+Take the bin plot on the left for example, the original activation values (blue line) ranges from [-10, 34] but with over 99% values between [-0.6, 1.7].
+More visualizations can be found in Figure 6.
+This phenomenon poses difficulties in minimizing the clipping error and is unfriendly for effective quantization.
+Figure 2: Illustration of imbalanced activation distributions.
 In the activation distribution of the full-precision model, the majority of values cluster near zero with sporadic large values, presenting challenges for low-bit quantization.
-Our method reduces the large values to enable easier quantization for improved performance.
-Property : Accurate temporal information is critical for quantized diffusion model performance.
-As shown in Equation 1, diffusion model predictions are highly correlated with time step information.
+Our method reduces the significant values to enable easier quantization for improved performance.
+Moreover, we observe that these significant values are important for generation performance preservation (see Figure 7).
+With the significant values being important and the small values appearing frequently, neither of them are negligible and need to be carefully quantized at the same time.
+Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped significant values, generating corrupted images.
+This inspires us to reshape the activation distributions via model finetuning to attain more quantization-friendly distributions, as depicted in Figure 1(a).
+Table 2: Time embedding type comparison.
+Bit-width Time Embedding FID $\downarrow$ sFID $\downarrow$ W8A8 PTQ Quantized 7.58 22.07 Full Precision 6.77 22.03 Ours 5.61 21.22 W4A8 PTQ Quantized 8.59 22.74 Full Precision 7.55 21.69 Ours 6.95 23.17
+Property : Accurate temporal information is critical for quantization performance.
+Diffusion model predictions are highly correlated with time step information.
 Specifically, time steps are transformed and projected into time embeddings and then added onto the image features, guiding the model to denoise correctly.
-Table 2 shows the performance comparison of quantizing time embeddings or preserving their full-precision values in the LDM-4 of LSUN-Bedrooms (Yu et al., 2015).
+Table 2 shows the performance comparison of quantizing time embeddings or preserving their full-precision values in the LDM-4 of LSUN-Bedrooms [39].
 Under W8A8 and W4A8, quantizing the time embeddings leads to an increase of 0.81 and 1.04 (relatively 15%) in FID, respectively.
@@
 Thus, preserving diffusion model performance under quantization requires accurate temporal information.
+Figure 3: Effect of decreasing different activations’ bit-width on the model performance.
+The generation failure of FeedForward layers emerges at 6 bits, while all other linear layers barely fail at 4 bits and all convolution layers only fail at 4 bits.
 Property : Different activations have different sensitivity towards decreasing bit-width.
 Low-bit activation quantization is often more difficult than weight quantization and remains the key challenge for diffusion model quantization.
-Compared with weights than can be categorized into linear layers and convolutional layers, activation types are more diverse and complicated.
-We hence examine the effects on model performance when quantizing different activations.
-Figure 3 shows how sensitive different activations are to quantization in Stable Diffusion.
-We quantize three different types of activations to lower bits while maintaining the others’ bit-width (8-bit).
-We observe that the FeedForward layer (Feng et al., 2023) activations cause generation failure at 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits.
-This phenomenon indicates that a small amount of activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.
-Table 2: FP & Quantized Time Embedding Comparisons.
-Settings FID $\downarrow$ sFID $\downarrow$
-W8A8 (Quantized Time Embedding) 7.58 22.07
-W8A8 (FP Time Embedding) 6.77 22.03
-W8A8 (TAQuant+$\text{SLA}_{\text{TE}}$) 5.61 21.22
-W4A8 (Quantized Time Embedding) 8.59 22.74
-W4A8 (FP Time Embedding) 7.55 21.69
-W4A8 (TAQuant+$\text{SLA}_{\text{TE}}$) 6.95 23.17
+Compared with weights that can be categorized into linear layers and convolutional layers, activation types are more diverse and complicated.
+Figure 3 shows the sensitivity of different layers’ activations to quantization.
+Specifically, we quantize three different types of activations to lower bits while maintaining the others’ bit-width to 8-bit, and observe how the decreasing bit-width effects generation performance.
+Surprisingly, we observe that the FeedForward layer [8] activations cause generation failure at as large as 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits.
+This indicates that a small amount of activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.
 2.3 Quantization via Efficient Selective Finetuning
-In this section, we first theoretically discuss the potential reasons for the failure of previous methods and the underlying rationale for weight fine-tuning.
+We first theoretically discuss the potential reasons for the failure of previous methods and the underlying rationale for weight fine-tuning.
 Then we introduce a finetuning method QuEST for diffusion models that can significantly boost low-bit performance with less time and memory usage.
-Figure 1 illustrates the proposed framework, which is a distillation-based finetuning strategy consisting of selective weight optimization and network-wise scaling factor optimization.
-To further address the loss of temporal information due to quantization, we propose an efficient time-aware activation quantizer.
+Figure 1(c) shows our framework, being a distillation-based finetuning strategy consisting of selective weight optimization and network-wise scaling factor optimization.
+To further address the loss of temporal information due to quantization, an efficient time-aware activation quantizer is proposed.
 2.3.1 Sufficiency of Finetuning
-We first review the underlying theoretical foundation underpinning previous methods, which typically employ the reconstruction-based post-training quantization approach.
+Motivated by the phenomenon that current PTQ methods completely fail when applied to low-bit diffusion models and driven by the phenomenon that the performance degradation caused by activation bit-width decrease is much more significant than weight bit-width (8-bit to 4-bit) in diffusion models, we would like to explore theoretically why previous PTQ methods fail and how to overcome this.
+We first review the underlying theory underpinning previous methods, which typically employ the reconstruction-based post-training quantization approach.
 Denote the diffusion model’s activations at time $t$ as $\mathbf{z}_{t}=[z_{1,t},z_{2,t},...,z_{n,t}]$, the task loss as $L(z_{n,t},\overline{z}_{n,t})$, where $n$ is the number of layers and $\overline{z}_{n,t}$ is the ground-truth.
@@
 We treat quantization as a type of perturbation and formulate the influence of activation quantization using Taylor expansion, assuming model weight $\mathbf{w}$ is frozen:
-$\mathbb{E}[L(z_{n,t}$ $+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]$
-$\approx\Delta^{T}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (4)
+$\mathbb{E}[L(z_{n,t}$ $+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]\approx\Delta^{T}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (3)
 where $\Delta$ is the activation perturbation, $\overline{\mathbf{g}}^{(\mathbf{z})}$ is the gradient and $\overline{\mathbf{H}}^{(z_{n,t})}$ is the Hessian matrix.
-According to (Li et al., 2021; Yuan et al., 2022), for a well-trained model, $\overline{\mathbf{g}}^{(z_{n,t})}=\nabla_{z_{n,t}}L$ approaches 00.
+According to [22, 40], for a well-trained model, $\overline{\mathbf{g}}^{(z_{n,t})}=\nabla_{z_{n,t}}L$ approaches 00.
 Thus the above equation can be simplified to:
@@
 This leads to a rather straightforward conclusion: Quantization error can be reduced by reconstructing the model output.
-However, under low-bit settings, the reasoning from Equation 3.3.1 to Equation 5 is inaccurate, where the activation perturbation $\Delta$ is too large for meaningful Taylor expansion.
+However, under low-bit settings, the reasoning from Equation 3 to Equation 4 is inaccurate, where the activation perturbation $\Delta$ is too large for meaningful Taylor expansion.
 Thus we have the following proposition:
@@
 $\mathbb{E}[L(z_{n,t}+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]$
-[…]
-[…]
+$\approx$ $2\epsilon^{T}\sum_{i=1}^{K}(\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}-z_{\text{FP}})+\frac{1}{2}\sum_{i=1}^{K}(\tilde{z}_{n,t}^{i}-z_{\text{FP}})^{T}\overline{\mathbf{H}}^{(z_{n,t}+(i-1)\epsilon)}(\tilde{z}_{n,t}^{i}-z_{\text{FP}})$ (5)
 where $\mathbf{w}_{n}$ is the weight for layer $n$ and $\tilde{z}_{n,t}^{i}=\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}$, $\epsilon$ is a perturbation small enough for Taylor expansion, $K$ is a constant and $\Delta=K\epsilon$.
+Proof is provided in Appendix B.
 Theorem 2.2 indicates that: To minimize the quantization error, ideally we can finetune $\mathbf{w}_{n}$ so that for any $i$, the weights can fit to the data/activation $\tilde{z}_{n-1,t}^{i}+(i-1)\epsilon$ which the full-precision counterpart may not fit, making the model more robust to varying inputs.
@@
 We also find that the second term in Equation 2.2 can be minimized simultaneously with the first term, so that the total quantization error can be reduced via a common goal: Activation output alignment via finetuning $\mathbf{w}_{n}$.
-Figure 3: Effect of decreasing different activations’ bit-width on the model performance in Stable Diffusion.
-FeedForward layers experience generation failures at 6 bits, while all other linear layers barely fail at 4 bits and all convolution layers only fail at 4 bits.
 2.3.2 Quantization via Selective Finetuning
@@
 Data-Free Partial-network-wise Training.
-To alleviate the need of substantial training data, we propose to adopt the general setting of Post-Training Quantization and construct the calibration set in a data-free manner.
-By feeding sampled Gaussian noises $x_{T}$ into the full-precision model and sample iteratively over different time steps using Equation 1, we can obtain the calibration data needed for finetuning the quantized model.
-In practice, the number of samples generated for each time step is 256, indicating that we only have to inference the full-precision model a few times to obtain the needed number of calibration samples.
-Moreover, inspired by the fact that we can view the activation quantization parameters as additional model parameters which all contribute to the final output, we propose a partial-network-wise training strategy.
-Different from quantization methods using layer-wise or block-wise reconstruction (Shang et al., 2023; Li et al., 2023a) that binds quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors for each time step and only train partial weight parameters (e.g.
-7% of the parameters in LDM-4).
-As depicted in Figure 1, the majority of the model weights remain unchanged at time step $t$, and only a small subset of parameters (including layer weights and scaling factors) related to this time step are updated using the full-precision model’s outputs as ground-truth.
+To alleviate the need of substantial training data, we construct the calibration set in a data-free manner.
+By feeding sampled Gaussian noises $x_{T}$ into the full-precision model and sample iteratively over different time steps, we can obtain the calibration data needed for finetuning the quantized model.
+In practice, the number of samples generated for each time step is 128 or 256, indicating that we only have to inference the full-precision model a few times to obtain the needed number of calibration samples.
+As depicted in Figure 1(c), to overcome the quantization difficulty demonstrated in property efficiently, we update partial model weights ($\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$) that only account for a small subset of parameters related to the time step $t$.
+The other weight parameters (such as $\mathbf{W}_{l^{{}^{\prime}}}$, $\mathbf{W}_{l^{{}^{\prime\prime}}}$) are frozen during optimization.
+For example on LDM-4 [32], at most 7% of the parameters are included for adjustment.
 The choices for the weights to be finetuned are discussed in the following sections.
-Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, our approach update the required parameters at once.
+To better conserve the performance and behavior of the full-precision model, we use both its intermediate activations and final outputs as the ground-truth for supervision.
+Figure 4 visualizes the generation results and demonstrates the effectiveness of this strategy in preserving image details.
+The activation quantization parameters can be viewed as additional model parameters, therefore we further propose a network-wise training strategy.
+Different from quantization methods using layer-wise or block-wise reconstruction [19, 34] that binds quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors for each time step together with the partial weight parameters.
+Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, we update the required parameters at once.
 In this way, we significantly save the time and memory needed for quantization.
 Note that the quantization parameters for the weights are not updated once initialized, in order to maintain memory efficiency during deployment.
-Table 3: Unconditional image generation results of LDM-4 models on LSUN-Bedrooms/Churches 256$\times$256. ”N/A” denotes generation failure. ”FP” denotes the full precision model.
-Dataset Method
-Bit-width
-(W/A)
-Size
+Efficient Time-aware Activation Quantization.
+As discussed in property , accurate time embedding quantization is critical.
+To discriminate between different time steps, we design a temporal quantizer specifically for activation quantization that adopts different sets of quantization parameters for different time steps.
+While these parameters consumes negligible memory and does not effect inference speed, learning independent parameters for each time step leads to immense time cost.
+Motivated by the fact that neighboring time steps yield similar functions in diffusion models [3], we cluster neighboring time steps so that they share the same quantization parameters and are calculated by the same calibration data $x_{t}$: $\mathbf{s}(t)=\mathbf{s}(t+1)=...=\mathbf{s}(t+\Delta t)$, where the time steps used for optimization are sampled uniformly, and each chosen time step’s statistics are used to represent $\Delta t$ adjacent ones.
+Though most time steps’ quantization parameters are approximated, this strategy saves the optimization time by $t$ times without compromising much performance.
+Section 3.2 further provides an analysis on the performance and efficiency trade-off.
+Table 3: Unconditional image generation results of LDM-4 models on LSUN-Bedrooms/Churches 256$\times$256. "N/A" denotes generation failure. "FP" denotes the full precision model.
 […]
-LSUN- Bedrooms (LDM-4) FP 32/32 1045.6 2.95
-PTQ4DM 8/8 279.1 4.75
-Q-Diffusion 8/8 279.1 4.53
-PTQ-D 8/8 279.1 3.75
-Ours 8/8 279.1 3.07
-PTQ4DM 4/8 148.4 -
-Q-Diffusion 4/8 148.4 5.37
-PTQ-D 4/8 148.4 5.94
-Ours 4/8 148.4 3.26
-PTQ4DM 4/4 148.4 N/A
-Q-Diffusion 4/4 148.4 N/A
-PTQ-D 4/4 148.4 N/A
-Ours 4/4 148.4 5.64
-LSUN- Churches (LDM-8) FP 32/32 1125.4 4.02
-PTQ4DM 8/8 330.6 5.54
-Q-Diffusion 8/8 330.6 6.94
-PTQ-D 8/8 330.6 6.40
-Ours 8/8 330.6 6.55
-PTQ4DM 4/8 189.9 -
-Q-Diffusion 4/8 189.9 7.80
-PTQ-D 4/8 189.9 7.33
-Ours 4/8 189.9 7.33
-PTQ4DM 4/4 189.9 N/A
-Q-Diffusion 4/4 189.9 N/A
-PTQ-D 4/4 189.9 N/A
-Ours 4/4 189.9 11.76
-Efficient Time-aware Activation Quantization.
-Temporal information is a unique and essential factor in diffusion models.
-As highlighted in (Shang et al., 2023; Li et al., 2023a), the variation of activation distribution over different time steps is fatal for conventional PTQ methods.
-Also, as discussed in Section 3.2, accurate time embedding quantization is critical.
-To discriminate between different time steps, we design the Time-aware Activation Quantizer (TAQuant) that adopts different sets of activation quantization parameters for different time steps.
-While these parameters consumes negligible memory and does not effect inference speed, learning independent parameters for each time step leads to immense time cost.
-Inspired by the fact that neighboring time steps yield similar functions in diffusion models (Choi et al., 2022), we use the simple strategy of clustering neighboring time steps so that they share the same calibration data and quantization parameters.
-Then the set of activation quantization parameters are formulated as:
-$\mathbf{s}=\{\mathbf{s}^{t}(x_{t}),\mathbf{s}^{2t}(x_{2t}),...,\mathbf{s}^{T}(x_{T})\},$ (7)
-where the time steps used for optimization are sampled uniformly, and each chosen time step’s statistics are used to represent $T/t$ adjacent ones.
-Though most time steps’ quantization parameters are approximated, this strategy saves the optimization time by $t$ times without compromising much performance.
-Section 4.3 further provides an analysis on the trade-off between time efficiency and model performance.
 Selective & Progressive Layer Alignment.
@@
 We progressively update these components due to their distinct, non-overlapping functionalities.
-Table 4: Class-conditional image generation results of the LDM-4 model on ImageNet 256$\times$256, containing 1,000 categories.
-Bit-width
-(W/A)
-Method
-Size
-[…]
-32/32 FP 1529.7 11.28 7.70 364.73
-8/8 Q-Diffusion 428.7 10.60 9.29 350.93
-PTQ-D 428.7 10.05 9.01 359.78
-Ours 428.7 10.43 6.07 365.12
-4/8 Q-Diffusion 237.5 9.29 9.29 336.80
-PTQ-D 237.5 8.74 7.98 344.72
-Ours 237.5 8.48 6.55 346.02
-4/4 Q-Diffusion 237.5 N/A N/A N/A
-PTQ-D 237.5 N/A N/A N/A
-Ours 237.5 5.98 7.93 202.45
-Table 5: Text-to-image generation results (512$\times$512) of Stable Diffusion v1.4 using COCO2014 validation set prompts.
-[…]
-32/32 FP 3279.1 31.50
-8/8 Q-Diffusion 949.0 31.43
-Ours 949.0 31.47
-4/8 Q-Diffusion 539.1 31.39
-Ours 539.1 31.50
-4/4 Q-Diffusion 539.1 N/A
-Ours 539.1 28.85
-Figure 4: Generated images compared with Q-Diffusion under two bit-width types: W8A8 and W4A4, using LDM-4 in LSUN-Bedrooms.
+Figure 4: Generated images compared with Q-Diffusion under two bit-width types: W8A8 and W4A4, using LDM-4 for LSUN-Bedrooms.
 Our method not only succeed well in generating low-bit images, but also aligns more closely with the full-precision counterpart.
-During a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations.
+In a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations.
 This implies that the time information operates independently from the primary network flow.
-Thus, we first refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}^{t}$ using:
-$\mathcal{L}_{\text{TE}}^{t}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||\overline{O}(e(t);\mathbf{w}_{l})-\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}^{t})||^{2}],$ (8)
-where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers, $e(t)$ is the fixed time embedding for time $t$, $\overline{O}(e(t);\mathbf{w}_{l})$ is the activation output of the full-precision model representing the ground-truth, and $\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}^{t})$ is the quantized output.
-This loss function indicates that the chosen weight parameters are consistently updated across different time steps to ensure robustness to diverse inputs, whereas only the subset of scaling factors relevant to each specific time step are adjusted for temporal information discrimination.
-Apart from time embedding layers, attention-related layers are also vital.
-As depicted in Figure 3, reducing activation bit-width of these layers degrades model performance the most dramatically.
-Denote $\mathbb{C}_{\text{A}}$ as the set containing all these layers and given the image calibration samples $x_{t}$, we optimize $\mathbf{w}_{l}$ and $\mathbf{s}^{t}$ but exclude the quantization parameters already optimized in Equation 8:
-$\mkern-18.0mu\mathcal{L}_{\text{A}}^{t}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||\overline{O}(x_{t};\mathbf{w}_{:l})-\tilde{O}(x_{t};\mathbf{w}_{:l},\mathbf{s}^{t}\setminus s_{\mathbb{C}_{\text{TE}}}^{t})||^{2}],\mkern-18.0mu$ (9)
+Thus, we first refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}(t)$:
+$\mathcal{L}_{\text{TE}}^{t}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||\overline{O}(e(t);\mathbf{w}_{l})-\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}(t))||^{2}],$ (6)
+where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers, $e(t)$ is the fixed time embedding for time $t$.
+$\overline{O}(e(t);\mathbf{w}_{l})$ is the final activation output of the full-precision model representing the ground-truth, and $\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}(t))$ is the quantized output.
+This objective function indicates that the chosen weight parameters are consistently updated across different time steps, so as to ensure robustness to diverse temporal inputs.
+Nevertheless, only the small subset of scaling factors relevant to each specific time step are adjusted for temporal information discrimination.
+Apart from time embedding layers, attention-related layers are also vital (according to property ).
+As depicted in Figure 3, reducing the activation bit-width of these layers degrades model performance the most dramatically.
+Denote $\mathbb{C}_{\text{A}}$ as the set containing all these layers and given the image calibration samples $x_{t}$, we optimize the corresponding weights $\mathbf{w}_{l}$ and quantization parameters $\mathbf{s}(t)$.
+However, we exclude the quantization parameters already optimized in Equation 6:
+$\mkern-18.0mu\mathcal{L}_{\text{A}}^{t}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||\overline{O}(x_{t};\mathbf{w}_{:l})-\tilde{O}(x_{t};\mathbf{w}_{:l},\mathbf{s}(t)\setminus s_{\mathbb{C}_{\text{TE}}}^{t})||^{2}],\mkern-18.0mu$ (7)
 where $\mathbf{w}_{:l}$ are the weight parameters of the first $l$ layers. This strategy resembles linear probing, where only the layer directly associated with the activation output is finetuned.
 Furthermore, we aim not only for each independent layer’s quantization error reduction, but also for the final generated images’ quality improvement.
-Therefore, the target task loss $\sum_{t}\mathcal{L}_{d}^{t}$ is introduced, where the ground-truth is approximated by the output of the full-precision counterpart.
+Therefore, we adopt the Task Loss Guidance strategy where the target task loss $\sum_{t}\mathcal{L}_{d}^{t}$ is introduced and the ground-truth is approximated by the output of the full-precision counterpart.
 By integrating Equation 6 and Equation 7, the final objective is formulated as follows:
 $\mathcal{L}$ $=\sum_{t}((\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{d}^{t})+(\mathcal{L}_{\text{A}}^{t}+\mathcal{L}_{d}^{t}))$
-$=\sum_{t}(\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{\text{A}}^{t}+2\mathbb{E}[||\overline{O}(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s}^{t})||^{2}]).$ (10)
+$=\sum_{t}(\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{\text{A}}^{t}+2\mathbb{E}[||\overline{O}(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s}(t))||^{2}]).$ (8)
 3 Experiments
 3.1 Experiment Settings
-To verify the effectiveness of our proposed method, we conduct experiments on three types of generation tasks: Unconditional image generation on LSUN-Bedrooms and LSUN-Churches datasets (Yu et al., 2015), class-conditional image generation on ImageNet (Deng et al., 2009), and text-to-image generation.
-The model architectures we quantize include LDM and Stable Diffusion (Rombach et al., 2021), and uses ”WnAm” to represent the quantization setting: n-bit weight quantization and m-bit activation quantization.
-DDIM samplers (Ho et al., 2020) are adopted for LDMs and the PLMS sampler (Liu et al., 2022) is used for Stable Diffusion.
+To verify the effectiveness of our proposed method, we conduct experiments on three types of generation tasks: Unconditional image generation on LSUN-Bedrooms and LSUN-Churches datasets [39], class-conditional image generation on ImageNet [5], and text-to-image generation.
+The model architectures we quantize include LDM and Stable Diffusion [32], and uses "WnAm" to represent the quantization setting: n-bit weight quantization and m-bit activation quantization.
+DDIM samplers [15] are adopted for LDMs and the PLMS sampler [26] is used for Stable Diffusion.
 We generate 256 samples per time step for constructing the calibration set, and train 20 epochs over each time step.
-The Adam optimizer (Kingma & Ba, 2017) is adopted and the learning rate for weight finetuning and scaling factor finetuning are set as $1e^{-5}$ and $1e^{-4}$ respectively.
+The Adam optimizer [18] is adopted and the learning rate for weight finetuning and scaling factor finetuning are set as $1e^{-5}$ and $1e^{-4}$ respectively.
+Table 4: Class-conditional generation results using LDM-4 on ImageNet 256$\times$256.
 […]
-4.2 Experiment Results
+[…]
+3.2 Experiment Results and Analysis
 Unconditional Generation: We evaluate the performance of our method over LDM-4 (LSUN-Bedrooms 256$\times$256) and LDM-8 (LSUN-Churches 256$\times$256) using the DDIM sampler with 200 and 500 time steps, respectively.
 Results are shown in table 3.
-We compare with methods including PTQ4DM (Shang et al., 2023), Q-Diffusion (Li et al., 2023a) and PTQ-D (He et al., 2023b).
-The results are evaluated using FID, where the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet (Li et al., 2023a).
-Table 6: Different component comparison and efficiency comparison on LDM-4 (LSUN-Bedrooms 256 $\times$ 256).
+The results are evaluated using FID, where the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet [19].
+Table 5: Text-to-image generation on Stable Diffusion (512$\times$512) using COCO2014 prompts.
+[…]
+Class-conditional Generation: We evaluate the performance using LDM-4 on ImageNet 256$\times$256 using the DDIM sampler (20 steps).
+As shown in Table 4, three metrics are used for evaluation.
+Note that sFID uses additional intermediate spatial features for calculation compared with FID.
+We can also see that FID is not a valid metric for ImageNet LDM-4 evaluation: All methods have lower FID when quantized to lower-bits, conflicting with human perception.
+We show that our method not only succeeds in W4A4 quantization, but also significantly improve the generation quality on higher bit settings.
+Under W8A8, our method is able to achieve 2.82 improvement over sFID and 8.13 increase in Inception Score.
+Under W4A8, we improve the SOTA PTQ method by 1.57 sFID and 8.77 IS.
+Text-to-image Generation: We use Stable Diffusion v1.4 as the model for quantization with the PLMS sampler sampling 50 time steps.
+Table 5 shows the results.
+Images are generated based on the 10,000 prompts sampled from the COCO2014 [24] validation set, and CLIP Score is calculated based on the ViT-B/16 backbone.
+Given the limited works done on Stable Diffusion, we can only compare with Q-Diffusion and the full-precision baseline.
+Table 6: Influence of task loss supervision on model performance.
+[…]
+Efficiency comparison and effects of components.
+Table 6 compares with and without the use of task loss $\mathcal{L}_{d}$.
+The results show that using the difference between the outputs of the quantized model and the full-precision counterpart for supervision is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for $\text{SLA}_{\text{TE}}$ and $\text{SLA}_{\text{A}}$, respectively.
+However, when the learning process is only supervised by the task loss, the performance degrades by 7.13 FID and 9.39 sFID for $\text{SLA}_{\text{TE}}$, indicating that the task loss alone is insufficient.
+Table 7 compares efficiency and performance between our method and the post-training quantization approach on the LSUN-Bedrooms dataset when using LDM-4.
+Our approach, while utilizing the same quantity of calibration data as the PTQ method, is more time-efficient and incurs only a 20% increase in GPU memory usage.
+We also show how each proposed component improves the generation performance.
+The results indicate that sequentially finetuning time embedding layers followed by finetuning attention-related layers leads to consistent performance improvements.
+Table 7: Component and efficiency comparisons on LDM-4 (LSUN-Bedrooms 256 $\times$ 256).
 Method
@@
 FP 32/32 - - - 1045.6 2.95
-PTQ (Li et al., 2023a) 4/4 5120 23.08 10334 148.4 428.8
-TAQuant 4/4 5120 11.52 9822 148.4 12.56
-+ $\text{SLA}_{\text{TE}}$ 4/4 5120 13.13 12134 148.4 6.41
-+ $\text{SLA}_{\text{A}}$ 4/4 5120 15.25 12450 148.4 5.78
-Class-conditional Generation: We evaluate the performance using LDM-4 on ImageNet 256$\times$256 using the DDIM sampler (20 steps).
-As shown in Table 4, three metrics are used for evaluation.
-Note that sFID uses additional intermediate spatial features for calculation compared with FID.
-We can also see that FID is not a valid metric for ImageNet LDM-4 evaluation: All methods have lower FID when quantized to lower-bits, conflicting with human perception.
-We show that our method not only succeeds in W4A4 quantization, but also significantly improve the generation quality on higher bit settings.
-Under W8A8, our method is able to achieve 2.82 improvement over sFID and 8.13 increase in Inception Score.
-Under W4A8, we improve the SOTA PTQ method by 1.57 sFID and 8.77 IS.
-Text-to-image Generation: We use Stable Diffusion v1.4 as the model for quantization with the PLMS sampler sampling 50 time steps.
-Table 5 shows the results.
-Images are generated based on the 10,000 prompts sampled from the COCO2014 (Lin et al., 2015) validation set, and CLIP Score is calculated based on the ViT-B/16 backbone.
-Given the limited works done on Stable Diffusion, we can only compare with Q-Diffusion and the full-precision baseline.
-4.3 Experimental Analysis
-Efficiency comparison and effects of different component.
-In Table 6, we present a comparison of efficiency between our method and the post-training quantization approach on the LSUN-Bedrooms dataset.
-Our approach, while utilizing the same quantity of calibration data as the PTQ method, is more time-efficient and incurs only a 20% increase in memory usage.
-We also show how each proposed component improves the generation performance.
-The results indicate that sequentially finetuning time embedding layers followed by attention-related layers leads to consistent improvements in model performance.
-Table 7 presents the comparison of with and without the use of task loss $\mathcal{L}_{d}$.
-The results show that using the difference between the outputs of the quantized model and the full-precision counterpart for supervision is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for $\text{SLA}_{\text{TE}}$ and $\text{SLA}_{\text{A}}$, respectively.
-However, when the learning process is only supervised by the task loss, the performance degrades by 7.13 FID and 9.39 sFID for $\text{SLA}_{\text{TE}}$, indicating that the task loss alone is insufficient.
-Table 7: Influence of task loss supervision on model performance.
-Method
-$\text{SLA}_{\text{TE}}$
-w/o $\mathcal{L}_{d}$
-$\text{SLA}_{\text{TE}}$
-w/ $\mathcal{L}_{d}$
-$\text{SLA}_{\text{A}}$
-w/o $\mathcal{L}_{d}$
-$\text{SLA}_{\text{A}}$
-w/ $\mathcal{L}_{d}$
-[…]
-[…]
+PTQ [19] 4/8 5120 23.08 10334 148.4 5.37
+Baseline 4/8 5120 11.52 9822 148.4 6.95
++ $\text{SLA}_{\text{A}}$ 4/8 5120 13.13 12134 148.4 4.41
++ $\text{SLA}_{\text{A}}$ + $\text{SLA}_{\text{TE}}$ 4/8 5120 15.25 12450 148.4 3.26
+Figure 5: Effect of uniformly sampled time steps number. Circle size represents quantization computation cost. Red square indicates the value we choose.
 Time step sampling frequency analysis.
 While time steps are sampled uniformly, will the performance consistently improve when we sample more frequently?
-Table 8 illustrates the trade-off between the increased number of sampled time steps (’T-Num’) and generation performance.
+Figure 5 illustrates the trade-off between the increased number of sampled time steps (’T-Num’) and generation performance.
 Metrics are calculated using 10,000 samples generated from LDM-4 on LSUN-Bedrooms.
@@
 Given the linearly increasing time cost associated with using more time steps, we opt for only 20 out of 200 time steps to enhance quantization efficiency.
-Table 8: Effect of number of uniformly sampled time steps for W4A4 LDM-4 on LSUN-Bedrooms dataset.
-T-Num 5 10 20 25 40 50
-FID $\downarrow$ 19.34 15.27 12.56 12.16 11.04 10.54
-sFID $\downarrow$ 54.00 37.47 28.78 27.99 26.42 25.86
 How QuEST influences model properties.
@@
 This indicates a higher concentration of values around the mean, which can effectively reduce both rounding and clipping errors.
+Integration with LoRA finetuning.
+Finetuning quantized models is not trivial.
+Thus, we use QALoRA [11] that extends LoRA [16] and finetune on the ImageNet 256$\times$256 dataset.
+The weight rank of 32 is used for the LoRA modules and the parameters are trained over 100 time steps for 160 epochs.
+We find that integrating the QALoRA technique leads to a 5.62 decrease in FID, indicating that finetuning the original layers is a better solution for performance preservation.
 4 Conclusion
-In this work, we propose QuEST, an efficient data-free finetuning framework for low-bit diffusion model quantization.
-Our method is motivated by the three underlying properties found in quantized diffusion models.
+We have proposed QuEST, an efficient data-free finetuning framework for low-bit diffusion model quantization.
+Our method is motivated and guided by the three underlying properties found in quantized diffusion models.
 We also theoretically demonstrate the sufficiency of finetuning, interpreting it as a way to enhance model robustness against large activation perturbations.
-To alleviate the performance degradation, we propose to finetune the time embedding layers and attention-related layers under the supervision of the full-precision counterpart.
-A time-aware activation quantizer is also introduced to handle varying time steps.
-Experimental results over three high-resolution image generation tasks demonstrate the effectiveness and efficiency of QuEST, achieving low-bit compatibility with less time and memory cost.
+To alleviate the performance degradation, we propose to finetune the time embedding layers and the attention-related layers under the supervision of the full-precision counterpart.
+Experimental results on three high-resolution image generation tasks (including Stable Diffusion) demonstrate the effectiveness and efficiency of QuEST, achieving low-bit compatibility with less time and memory cost.
+Appendix A Related Works
+A.1 Diffusion Model Inference
+Diffusion models [15, 32] generate samples via a learned iterative denoising process.
+During inference, the initial input is sampled from a Gaussian distribution: $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the final output $x_{0}$ is obtained by removing noises from $x_{T}$ subsequently using:
+$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\tilde{\boldsymbol{\mu}}_{\theta,t}(x_{t}),\tilde{\beta_{t}}\mathbf{I}).$ (9)
+In this Markov Chain, $\tilde{\boldsymbol{\mu}}_{\theta,t}$ and $\tilde{\beta_{t}}$ are calculated from the model’s output.
+For a typical diffusion model, the denoising stage requires hundreds to thousands of iterations.
+This laborious sequential process makes efficient inference and high-resolution generation extremely challenging.
+Latent Diffusion Models (LDMs) [32] were developed to overcome the challenges.
+They encode the data into a latent space and then apply the diffusion process within this space, enabling high-resolution image generation with fewer time steps.
+The output images are produced by decoding the latent representations.
+LDM’s concept and network quantization complement each other, mutually enhancing the efficiency of diffusion models.
+Practically, diffusion models typically adopt a UNet architecture [33], with some modules substituted by attention blocks.
+In a latent diffusion model, an encoder and a decoder are incorporated for projecting images into latent representations and subsequently generating images from the latent space, respectively.
+Usually encoders and decoders are lightweight and computationally inexpensive, so our focus is on quantizing the UNets in LDMs, in alignment with other works (as described below).
+A.2 Diffusion Model Quantization
+Model quantization is a dominant technique for optimizing the inference memory and speed of deep learning models by reducing the precision of the tensors used in computation.
+Few researches were done for quantizing diffusion models and they fall into two mainstream frameworks: Quantization-Aware Training (QAT) [17, 20, 21, 38] and Post-Training Quantization (PTQ) [22, 25, 28, 36, 37].
+QAT methods such as Q-DM [20] consider value precision as an additional optimization constraint, training all parameters from scratch with the target task objective.
+While it is effective for low-bit quantization, it is even more resource-intensive than training a full-precision model.
+Efficient-DM [11] further uses Low-rank adapters (LoRA) [16] to reduce training cost.
+However, it introduces extra weight parameters and still requires substantial training iterations.
+Post-Training Quantization methods aim to calibrate the quantization parameters based on a small calibration set.
+Pioneering works such as PTQ4DM [34] and Q-Diffusion [19] focus on sampling the outputs of the full-precision model under different time steps with certain probabilities to construct the calibration set.
+ADP-DM [35] partition time steps into different groups and optimize each group with differentiable search algorithms.
+PTQ-D [12] decomposes the quantization noise and correct them individually with full-precision model statistics.
+However, these methods generally fail under 4-bit or lower bit-width.
+To enable low-bit compatibility with high efficiency, our method adopts the general framework of PTQ but further introduces an efficient weight fine-tuning strategy.
+Appendix B Proof for Theorem 2.2
+We provide the detailed proof for Theorem 2.2 here. The notations are consistent with the ones in the main paper.
+Since the perturbation $\Delta$ is too large for accurate Taylor expansion, we can resolve it by introducing a new perturbation $\epsilon=\Delta/K$, where we divide $\Delta$ by a constant $K$ so that $\epsilon$ is small enough for approximation.
+Then, Equation 3 is rewritten as follows:
+$\mathbb{E}[L(z_{n,t}+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]$
+$=$ $\sum_{i=1}^{K}\left(\mathbb{E}[L(z_{n,t}+\frac{i}{K}\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t}+\frac{i-1}{K}\Delta;\mathbf{w})]\right)$
+$\approx$ $\sum_{i=1}^{K}\left(\epsilon^{T}\overline{\mathbf{g}}^{(z_{n,t}+(i-1)\epsilon)}+\frac{1}{2}\epsilon^{T}\overline{\mathbf{H}}^{(z_{n,t}+(i-1)\epsilon)}\epsilon\right),$ (10)
+where the approximation step follows Taylor expansion and only the first two main components are kept.
+The first term in Equation B cannot be ignored because samples such as $z_{n,t}+(i-1)\epsilon$ may not be included in the learned distribution of the model.
+The second term can still be minimized by reconstruction since only the difference between quantized model output and ground-truth matters.
+In the following, we temporarily exclude the second term for simplicity since it can always be minimized through aligning the activation outputs.
+(a) Activation Distribution on Conditional LDM4 (ImageNet 256 $\times$ 256)
+(b) Activation Distribution on Unconditional LDM4 (LSUN-Bedrooms 256 $\times$ 256)
+Figure 6: Illustrations of imbalanced activation distributions on conditional LDM4 (ImageNet 256$\times$256) and unconditional LDM4 (LSUN-Bedrooms 256$\times$256).
+Given the objective function (MSE loss) of diffusion models, we analyze that:
+$\sum_{i=1}^{K}\epsilon^{T}\overline{\mathbf{g}}^{(z_{n,t}+(i-1)\epsilon)}$ $=2\epsilon^{T}\sum_{i=1}^{K}(\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}-\overline{z}_{n,t})$
+$\approx$ $2\epsilon^{T}\sum_{i=1}^{K}(\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}-z_{\text{FP}}),$ (11)
+where $\mathbf{w}_{n}$ is the weight for layer $n$, $\tilde{z}_{n-1,t}^{i}$ is the activation of the $(n-1)$th layer in a quantized model to get $z_{n,t}+(i-1)\epsilon$.
+Ground-truth $\overline{z}_{n,t}$ can be approximated by the full-precision output $z_{\text{FP}}$.
+We see that $\tilde{z}_{n-1,t}^{i}$ and $z_{\text{FP}}$ cannot be changed, thus to minimize Equation B, we need to finetune $\textbf{w}_{n}$.
+From a general perspective, Equation B also indicates that the model has not converged well to a local minimum given the perturbed inputs, thus when we finetune the model layers given the quantized inputs, we are actually training the model towards convergence over new samples and increasing its robustness.
+Appendix C Examples of Imbalanced Activation Distributions
+Apart from Figure 2, we show that the imbalance in the activation distribution is a common phenomenon in different model structures and datasets.
+In Figure 6, we show more results of activation distributions of latent diffusion models on ImageNet 256 $\times$ 256 and LSUN-Bedrooms 256 $\times$ 256.
+Appendix D Importance of large values in activations
+As shown in Figure 2, quite a few values are rather large and diversely distributed.
+These values pose difficulties on activation quantization, and being rather important and not negligible.
+To demonstrate this, we corrupt certain tokens in the activation outputs of the diffusion model and check the corresponding generated images.
+The corruption is done by setting the token values as all zeros.
+As shown in Figure 7, we compare two settings: (1) corrupt a certain number of tokens randomly; (2) corrupt the same number of the tokens with the largest values.
+We see that when corrupting randomly, generation performance is hardly effected.
+However, corrupting the same amount of tokens (even only one token) with the largest values leads to significantly degenerated images.
+Figure 7: Comparison of different corruptions made on different tokens.
+Appendix E Generated image examples for different tasks
+E.1 Unconditional Image Generation
+The generated images for LDM-4 on LSUN-Bedrooms 256$\times$256 under different bit-widths are shown in Figure 8. Images for LSUN-Churches 256$\times$256 are shown in Figure 9.
+(a) Full Precision
+(b) W8A8
+(c) W4A8
+(d) W4A4
+Figure 8: Unconditional image generation examples for LSUN-Bedrooms 256$\times$256.
+(a) Full Precision
+(b) W8A8
+(c) W4A8
+(d) W4A4
+Figure 9: Unconditional image generation examples for LSUN-Churches 256$\times$256.
+E.2 Class-conditional image generation
+Figure 10 shows the generated images for 3 different classes.
+(a) Full Precision
+(b) W8A8
+(c) W4A8
+(d) W4A4
+Figure 10: Conditional image generation results for ImageNet 256$\times$256.
+Figure 11: Text-to-image generation results on Stable Diffusion.
+E.3 Text-to-image generation
+Figure 11 shows the generated images for two different prompts using Stable Diffusion v1.4 under different bit-width.
+Appendix F Limitations and Broader Impacts
+The primary objective of this paper is to further the research in enhancing the efficiency of diffusion models.
+While it confronts societal consequences akin to those faced by research on generative models, it is important to recognize the potential impacts that quantized models could have on current techniques, including watermarking and safety checking.
+Inappropriate integration of current methodologies may result in unforeseen performance issues, a factor that deserves attention and awareness.

### v4

@@
 Abstract
-The practical deployment of diffusion models still suffers from the high memory and time overhead.
-While quantization paves a way for compression and acceleration, existing methods unfortunately fail when the models are quantized to low-bits.
-In this paper, we empirically unravel three properties in quantized diffusion models that compromise the efficacy of current methods: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules.
-To alleviate the intensified low-bit quantization difficulty stemming from the distribution imbalance, we propose finetuning the quantized model to better adapt to the activation distribution.
-Building on this idea, we identify two critical types of quantized layers: those holding vital temporal information and those sensitive to reduced bit-width, and finetune them to mitigate performance degradation with efficiency.
-We empirically verify that our approach modifies the activation distribution and provides meaningful temporal information, facilitating easier and more accurate quantization.
-Our method is evaluated over three high-resolution image generation tasks and achieves state-of-the-art performance under various bit-width settings, as well as being the first method to generate readable images on full 4-bit (i.e.
-W4A4) Stable Diffusion.
-Code is available here.
+The practical deployment of diffusion models is still hindered by the high memory and computational overhead.
+Although quantization paves a way for model compression and acceleration, existing methods face challenges in achieving low-bit quantization efficiently.
+In this paper, we identify imbalanced activation distributions as a primary source of quantization difficulty, and propose to adjust these distributions through weight finetuning to be more quantization-friendly.
+We provide both theoretical and empirical evidence supporting finetuning as a practical and reliable solution.
+Building on this approach, we further distinguish two critical types of quantized layers: those responsible for retaining essential temporal information and those particularly sensitive to bit-width reduction.
+By selectively finetuning these layers under both local and global supervision, we mitigate performance degradation while enhancing quantization efficiency.
+Our method demonstrates its efficacy across three high-resolution image generation tasks, obtaining state-of-the-art performance across multiple bit-width settings.
+Code is available at https://github.com/hatchetProject/QuEST.
 1 Introduction
-Diffusion models [6, 15, 32] have achieved great success in the field of image generation recently.
-Given the high capacity and flexibility of diffusion models, there still remains two challenges that limit their efficiency [4].
-The first obstacle is the time-consuming denoising process which requires hundreds to thousands of inference time steps that slows down the generation speed drastically.
-The other one is the increasing model size driven by the need of better image fidelity and higher image resolutions.
+Diffusion models [6, 13, 28] have recently achieved remarkable success in image generation.
+However, this success comes at the cost of two major obstacles that limit their efficiency [3].
+The first obstacle is the denoising process which requires hundreds to thousands of inference time steps, slowing down the generation speed drastically.
+The other is the increasing model size, driven by demands for better image fidelity and higher image resolutions.
 Both factors contribute to considerable latency and increased computational requirements, impeding the application of diffusion models to real-world settings where both time and computational power are carefully restricted.
-Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously [9].
+Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously [8], making it a natural solution for deploying diffusion models efficiently.
 It aims to compress high-bit model parameters into low-bit approximations with negligible performance degradation.
-For example, 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and 8$\times$ memory reduction theoretically [23].
-The benefits will increase consistently when the bit-widths are further lowered.
+For example, 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and memory reduction theoretically [22].
 Hence, low-bit quantization of diffusion models emerges as a viable approach for efficiency enhancement.
-However, existing diffusion model quantization approaches fail under low-bit settings.
-These methods focus either on timestep-aware calibration data construction [19, 34] or quantization noise correction [12], targeting at tailoring existing quantization techniques to the properties of diffusion models that are distinct from other model types such as CNNs [31] and ViTs [2].
-Such approaches overlook the intrinsic diffusion model mechanisms that are related with quantization, leading to inconsistencies between method deployment and model characteristics.
-Therefore, we intend to identify unique features in diffusion models that are associated with quantization.
-In this paper, we reveal three properties of quantized diffusion models that hinder effective quantization: The activation distributions are likely to be imbalanced, where the majority of values are close to 0 but the important and large values are rare and appear inconsistently.
-Conventional quantization methods tend to either approximate the large and sparse values, inadequately estimating the numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error.
-Diffusion models exhibit varying functions at distinct time steps [3], such as forming object outlines at early time steps and constructing details latter.
-Thus, preserving accurate temporal information is important during quantization.
-However, current methods struggle to represent extensive time steps with limited bit-width, leading to similar time embeddings across different time steps and results in error accumulation.
-Diffusion models typically have complex network architectures, incorporating convolutional layers and linear layers that construct various kinds of attention blocks [7] and residual blocks [10].
-Whereas previous works consider each model component as equally important and apply quantization uniformly, we reveal that certain layers are particularly sensitive to perturbations from quantization, while others are more resilient.
-Table 1: Comparison with different frameworks.
+Unfortunately, existing diffusion model quantization methods that perform well at higher bit-widths face significant limitations in low-bit settings: some only adjust the quantization parameters and fail under low-bit conditions [18, 30, 10], while others succeed but require substantial computational resources comparable to training a diffusion model from scratch [21, 32].
+In this work, we aim for efficient low-bit quantization, thereby circumventing the latter choice of resource-intensive training.
 Method Data Time & Memory Low-bit Fully Free Efficient Compatible Quantized PTQ [18] QAT [21] EfficientDM [9] Ours
-Addressing the revealed properties, we propose a novel quantization approach for diffusion models, termed QuEST (Quantization via Efficient Selective FineTuning).
-Confronting property , we first theoretically justify that weight finetuning can enhance model robustness towards large perturbations in activations under low-bit settings for reducing quantization error, whereas previous methods struggle to find appropriate quantization parameters.
-Then we empirically show that by finetuning the model, the activation distributions are actually modified to be more amenable to quantization by reducing the significant values (large yet sparse), as shown in Figure 1(a).
-Following the idea of weight finetuning, we compare different bit-width models (Figure 1(b)) and identify two types of layers that are responsible for performance degradation: Time embedding layers exhibiting property and attention-related layers corresponding to property .
-Thus, we selectively and progressively finetune these small layer subsets in conjunction with all activation quantization parameters, as illustrated in Figure 1(c).
-With less than 10% of the total parameters involved, QuEST not only substantially enhances low-bit quantized model performance, but is also notably time-efficient and requires only supervision from the full-precision model in a data-free manner.
+Table 1: Comparison with different frameworks. Our method is both efficient and effective for low-bit diffusion model quantization, also achieving a reduced overall bit-width.
+We first reveal the current challenge within diffusion models that impede the effectiveness of current efficient low-bit quantization methods [18, 31, 15].
+As illustrated in Fig.
+1(a): activation distributions tend to be imbalanced, with most values clustering near zero, while essential high-magnitude values are sparse and inconsistently distributed.
+Existing quantization methods [30, 18, 19] either approximate large and sparse values, inadequately estimating numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error.
+To overcome this challenge, we propose to adjust the activation distributions via weight finetuning, where its feasibility is justified both theoretically and empirically.
+Nevertheless, finetuning the entire diffusion model is a highly computationally-expensive and time-consuming process, requiring over 80GB memory and numerous hours [5, 21].
+Thus, developing an efficient finetuning strategy tailored for diffusion model quantization is important.
+To facilitate efficient quantization, we further identify two key properties of quantized diffusion models that unlock new opportunities: diffusion models exhibit varying functions at distinct time steps [2], therefore preserving accurate temporal information is important during quantization; and diffusion models possess complex network architectures, incorporating various types of modules.
+Whereas previous works consider each module as equally important and apply quantization uniformly, we reveal that certain modules are particularly sensitive to perturbations from quantization, while others are more resilient.
+Figure 1: Overview of our observations and method.
+(a) Illustration of the challenge we identified in low-bit diffusion model quantization and a potential solution.
+We propose to ease the quantization difficulty by refining the activation distribution to make it more quantization-friendly.
+(b) Overview of property and property , whose importance are identified based on their impact on the model’s generation performance, making them suitable module candidates for efficient finetuning.
+(c) Framework of the proposed method.
+$\textbf{W}_{\text{TE}},\textbf{W}_{\text{A}},\textbf{W}_{\text{F}}$ are the weights of the time embeddings layers, attention-related layers, and other frozen layers, respectively.
+s is the quantization parameter.
+To train with efficiency, we adopt a selective and progressive finetuning strategy, incorporating temporal layer alignment (TLA) and critical module alignment (CMA).
+A global loss is also used for network-level guidance, improving the generated image quality.
+Based on the above findings, we propose a novel quantization approach for diffusion models, termed QuEST (Quantization via Efficient Selective FineTuning).
+Confronting the revealed quantization challenge, we first theoretically justify that weight finetuning can enhance model robustness toward large activation perturbations in low-bit settings, thereby reducing quantization error.
+In contrast, previous methods have struggled to properly balance clipping error and rounding error.
+Then we empirically demonstrate that by finetuning the model weights, the activation distributions are modified to be more amenable to quantization.
+As shown in Fig.
+1(a), the activation distribution is adjusted by reducing the amount of large, sparse values and enhancing the compactness of the value distribution.
+Following the idea of weight finetuning, we compare the effects of quantizing different modules of diffusion models (Fig.
+1(b)) and identify two types of layers as primary culprits to performance degradation: time embedding layers exhibiting property and attention-related layers associated with property .
+Consequently, we selectively and progressively finetune the small subsets of identified layers in conjunction with all activation quantization parameters, as illustrated in Fig.
+1(c).
+The learning objective is crafted to align the quantized model with its full-precision counterpart at both local and global levels.
+Involving less than 7% of the total parameters, QuEST not only substantially enhances low-bit quantized model performance, but is also notably time-efficient and can be conducted in a data-free manner.
 Our contributions are summarized as follows:
-  • We uncover and validate three properties that present challenges to low-bit diffusion model quantization: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules.
-We also theoretically discuss why previous methods are expected to fail and justify the sufficiency of weight finetuning under low-bit settings.
-  • Addressing the identified properties, we introduce QuEST, a new finetuning strategy that trains the model selectively and progressively, achieving low-bit quantization capability with high time and memory efficiency.
-An efficient activation quantizer is also proposed for time information estimation.
-Our method can be executed in a data-free and parameter-efficient manner.
+  • We identify the current challenge in low-bit diffusion model quantization that hinders effective low-bit quantization, and propose to adjust the activation distributions via weight finetuning for easier quantization.
+Both theoretical and empirical discussions are provided to justify the feasibility of this strategy under low-bit settings.
+  • We uncover and validate two properties in quantized diffusion models as the main factors for degraded performance.
+Motivated by the identified properties, we introduce QuEST, a parameter-efficient finetuning strategy that trains the diffusion model selectively and progressively, achieving low-bit quantization capability with time and memory efficiency.
   • Experiments on three high-resolution image generation tasks over four models demonstrate the superiority of our method, achieving state-of-the-art performance under various bit-width settings.
-We are also the first method to generate reasonable images for full 4-bit Stable Diffusion.
-The detailed related works on diffusion model inference and quantization are referred to appendix for space restriction.
-Table 1 is a compact summary that shows the unique capabilities of our method, where we are able to overcome the limitations of previous quantization methods.
-In comparison with EfficientDM [11], we include a brief discussion here: Technically, EfficientDM aims to finetune via low-rank weight representations [16].
-However, it introduces additional parameters for each layer and requires extensive iterations of training on LoRA weights.
-Our method instead does not include additional parameters and only needs to finetune the model for limited iterations with the same batch size.
-We have also found that EfficientDM does not quantize the matrix multiplications in the attention mechanism, while we quantize the matrices to 8-bits.
-Moreover, EfficientDM always keeps the first three layers and the last layer as full 8-bit (weight and activation), as well as keeping skip-connections and some linear layers as full-precision during implementation.
-Our method instead only kept the first and last layer as 8-bit and quantize all other layers to the required bit-width.
-Figure 1: Overview of our findings and method.
-(a) Illustration of property and a way to mitigate the quantization difficulty.
-(b) Demonstration of how property and property are identified via comparing different bit-width models.
-(c) Framework of the proposed method at time step $t$.
-To address the quantization difficulty and achieve more quantization-friendly activation distribution efficiently, we reduce significance by selectively and progressively updating partial model layers’ weights.
-$e(t)$ and $x_{t}$ are the input time embedding and input image representation respectively.
-While $\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$ are selected layer weights to be optimized using Equation 2.3.2, $\mathbf{W}_{l^{{}^{\prime}}}$ and $\mathbf{W}_{l^{{}^{\prime\prime}}}$ are frozen during finetuning.
-More details are included in Section 2.3.2.
+2 Related Works
+2.1 Diffusion Model Inference
+Diffusion models [13, 28] generate samples via an iterative denoising process.
+During inference, the initial input is sampled from a Gaussian distribution: $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the final output $x_{0}$ is obtained through a denoising process:
+$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\tilde{\boldsymbol{\mu}}_{\theta,t}(x_{t}),\tilde{\beta_{t}}\mathbf{I}),$ (1)
+where $\tilde{\boldsymbol{\mu}}_{\theta,t}$ and $\tilde{\beta_{t}}$ are calculated from the model’s output.
+This denoising process in a typical diffusion model requires tens to thousands of iterations, making efficient inference extremely challenging.
+Practically, diffusion models typically adopt a UNet architecture [29], incorporating an encoder and a decoder.
+Usually, encoders and decoders are lightweight and computationally inexpensive, so our focus is on quantizing the UNet structures in latent diffusion models, in alignment with the other works.
+2.2 Diffusion Model Quantization
+Model quantization is a dominant technique for optimizing the inference memory and speed of deep learning models by reducing the precision of the tensors used in computation.
+The researches for diffusion model quantization fall into three categories: Quantization-Aware Training (QAT) [16, 21, 20], Post-Training Quantization (PTQ) [19, 24, 34, 30, 18, 35], and Parameter-Efficient Fine-Tuning methods [9].
+QAT methods [21] train all parameters from scratch, being effective for low-bit quantization but are extremely resource-intensive.
+PTQ methods [33, 10, 15, 36, 31] calculate the quantization parameters based on a small calibration set, offering better efficiency.
+However, PTQ methods often rely on complex designs and fail at lower bit-widths.
+To achieve low-bit compatibility with high efficiency, Parameter-Efficient Fine-Tuning methods were proposed.
+The representative work EfficientDM [9] trains a low-rank adapter (LoRA) [14] for each layer to reduce training costs, and successfully scales to W4A4.
+Our proposed method also adopts a parameter-efficient finetuning strategy, and differs from EfficientDM in the following aspects: Firstly, EfficientDM introduces extra weight parameters, requiring substantial training iterations on the LoRA weights.
+Our method instead does not include additional parameters.
+Secondly, EfficientDM does not quantize the matrix multiplications in the attention mechanism, as well as certain linear layers.
+Our method quantizes all layers, and is more time-efficient.
+Tab.
+1 summarizes the differences between our method with the other works.
 3 Methodology
@@
 The quantization and dequantization processes are performed on both model weights and layer outputs (also termed as ’activation’).
-Equation 1 indicates that the quantization error is composed of two factors: The clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship [22].
-While previous approaches strive for an optimal balance between the two errors, they neglect the intrinsic mechanisms in quantized diffusion models.
-In the next section, we investigate three properties of diffusion models that relate to quantization performance and show why typical quantization methods tend to fail.
-2.2 Quantization-aware Properties of Diffusion Models
-Property : The activation outputs are likely to have a majority of values close to zero, with the other values being numerically large and important.
-Previous works [19, 34] paid attention to the varying activation distribution over different time steps, but none of them focused on the activation distribution itself.
-In Figure 2, we cluster activation values into uniformly distributed bins and manipulate the frequency for better visualization.
-We find that in some layers, though the majority of values are close to 0, there exists values that are relatively large and diverse, which we term as "significant values" (circled in green).
-Take the bin plot on the left for example, the original activation values (blue line) ranges from [-10, 34] but with over 99% values between [-0.6, 1.7].
-More visualizations can be found in Figure 6.
+Eq.
+2 indicates that the quantization error is composed of two factors: the clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship [19].
+While previous approaches strive for an optimal balance between the two errors, they neglect the intrinsic characteristics in quantized diffusion models.
+In the following sections, we first examine the current challenges in diffusion model quantization and outline our finetuning motivation, providing theoretical justification.
+We then identify two properties that enable efficient finetuning, forming the basis of our proposed method.
+Figure 2: Illustration of the imbalanced activation distributions.
+In the full-precision model, the majority of values cluster near zero with sporadic large values, presenting challenges for low-bit quantization.
+Our method refines the activation distributions by eliminating the large and sparse values, enabling easier quantization.
+3.2 The Challenge in Low-bit Quantization
+Previous works [18, 15] primarily address the varying activation distributions across different time steps, facilitating diffusion model quantization at higher bit-widths.
+However, these methods experience failure in low-bit settings.
+To investigate the potential reason, we focus on the activation distribution itself and conduct a layer-wise analysis, revealing the following challenge in full-precision diffusion models that impedes effective quantization:
+Challenge: Despite the majority of values being close to zero in the activation outputs, there exist numerically large and sparse values holding significant importance.
+Fig.
+2 provides a detailed analysis, where activation values are clustered into uniformly distributed bins.
+We find that in some layers, though the majority of values are close to zero, there exist values that are relatively large and diverse (circled in green).
+Take the bin plot on the left as an example, the original activation values (blue line) range from [-10, 34] but with most values between [-0.6, 1.7].
+Visualizations for more layers can be found in Appendix 9.
 This phenomenon poses difficulties in minimizing the clipping error and is unfriendly for effective quantization.
-Figure 2: Illustration of imbalanced activation distributions.
-In the activation distribution of the full-precision model, the majority of values cluster near zero with sporadic large values, presenting challenges for low-bit quantization.
-Our method reduces the significant values to enable easier quantization for improved performance.
-Moreover, we observe that these significant values are important for generation performance preservation (see Figure 7).
-With the significant values being important and the small values appearing frequently, neither of them are negligible and need to be carefully quantized at the same time.
-Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped significant values, generating corrupted images.
-This inspires us to reshape the activation distributions via model finetuning to attain more quantization-friendly distributions, as depicted in Figure 1(a).
-Table 2: Time embedding type comparison.
-Bit-width Time Embedding FID $\downarrow$ sFID $\downarrow$ W8A8 PTQ Quantized 7.58 22.07 Full Precision 6.77 22.03 Ours 5.61 21.22 W4A8 PTQ Quantized 8.59 22.74 Full Precision 7.55 21.69 Ours 6.95 23.17
-Property : Accurate temporal information is critical for quantization performance.
-Diffusion model predictions are highly correlated with time step information.
-Specifically, time steps are transformed and projected into time embeddings and then added onto the image features, guiding the model to denoise correctly.
-Table 2 shows the performance comparison of quantizing time embeddings or preserving their full-precision values in the LDM-4 of LSUN-Bedrooms [39].
-Under W8A8 and W4A8, quantizing the time embeddings leads to an increase of 0.81 and 1.04 (relatively 15%) in FID, respectively.
-This indicates that inaccurate time embedding quantization can lead to large degradation in generation performance, yielding even larger corruptions under lower bit-widths.
-Due to the different focus diffusion models have on image generation at different time steps, we infer that inaccurate time embeddings can cause mismatched input image and model functionality, resulting in possible oscillations in the sequence of noise removal.
-Thus, preserving diffusion model performance under quantization requires accurate temporal information.
+Moreover, these large and sparse values are important for generation performance preservation.
+We find that when replacing the few tokens with maximum values by random noises, the generated images’ quality is critically degraded (as shown in Appendix 10).
+With these large values being important and the small values appearing frequently, neither of them is negligible and needs to be carefully quantized at the same time.
+Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped values, generating corrupted images.
+This inspires us to refine the activation distributions to attain more quantization-friendly distributions, as depicted in Fig.
+1(a).
+However, the activation distribution cannot be directly manipulated.
+To address this, we instead finetune the model weights under quantization constraints, producing a new yet similar full-precision model whose quantized counterpart maintains performance comparable to the original full-precision model.
+Our experiments interestingly reveal that the proposed finetuning strategy effectively eliminates the large and sparse values (Fig.
+2), reducing quantization difficulty.
+We detail our approach in Sec.
+3.3 and provide further theoretical analysis in Sec.
+3.4.
+3.3 Quantization via Efficient Selective Finetuning
+In this section, we introduce QuEST, an efficient finetuning method for diffusion models that can significantly boost low-bit performance with less time and memory usage.
+We also present the two unique properties in quantized diffusion models, which serve as the foundation for the design of our method.
+Fig.
+1(c) illustrates our approach.
+3.3.1 Data-free Efficient Network-wise Training.
+We first present the general training pipeline of our method.
+To alleviate the need for substantial training data, we construct the calibration set in a data-free manner.
+By feeding random Gaussian noises $x_{T}$ into the full-precision model and sampling over different time steps, we can obtain the calibration data needed for finetuning the quantized model.
+In practice, we only have to infer the full-precision model a few times to gather the needed number of calibration samples, totaling 128 or 256 samples per time step.
+As depicted in Fig.
+1(c), to overcome the quantization challenge efficiently, we update partial model weights ($\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$) that only account for a small subset of parameters related to the time step $t$.
+The remaining weight parameters $\mathbf{W}_{\text{F}}$ are kept frozen during optimization.
+We also fix the weight quantization parameters during training, reducing the amount of parameters that need to be optimized.
+For instance, in LDM-4 [28], no more than 7% of the parameters are adjusted.
+The choices for the weights to be finetuned will be discussed in the following sections.
+The activation quantization parameters can be viewed as additional model parameters.
+Therefore, we further propose a network-wise training strategy.
+Different from quantization methods using layer-wise or block-wise reconstruction [18, 30] that bind quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors together with the partial weight parameters.
+Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, we update the required parameters at once.
+In this way, we significantly save the time and memory needed for quantization.
+3.3.2 Temporal Layer Alignment
+The inference process of diffusion models is highly dependent on the temporal information.
+Specifically, integer time steps are transformed into time embeddings through one or two linear layers, then added to the intermediate model features.
+Motivated by this observation, we make the following analysis that is consistent with previous works [15, 31]:
+Property : Although time embeddings depend solely on time steps and are easily obtainable, precise temporal information is crucial for optimal quantization.
+TE Setting (W8A8 Model) FID $\downarrow$ sFID $\downarrow$ TE Setting (W4A8 Model) FID $\downarrow$ sFID $\downarrow$ PTQ 7.58 22.07 PTQ 8.59 22.74 FP 6.77 22.03 FP 7.55 21.69 Ours 5.61 21.22 Ours 6.95 23.17
+Table 2: Ablations on time embedding (TE) settings. Finetuning the TE layers with our method surpasses full-precision embeddings, while the latter outperforms standard quantized ones.
+Tab.
+2 provides an empirical justification, where we quantitatively show the performance drop when quantizing time embeddings to different bit-width.
+Under W8A8 and W4A8 bit-width settings, solely quantizing the time embeddings can lead to an increase of 0.81 and 1.04 (relatively 15%) in FID, respectively.
+We infer the reason is that inaccurate time embeddings can cause mismatched input and model functionality, resulting in possible oscillations in the sequence of noise removal.
+Previous works either propose to learn dynamic quantization parameters across different time steps through a simple network [31], or calibrate the time embedding layers and projection layers across all time steps [15].
+We instead focus on finetuning the time embedding layers, adjusting fewer modules without introducing additional parameters.
+The results in Tab.
+2 also suggest that our method can improve the quantization performance, even surpassing the full-precision baseline.
+Concretely, in a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations.
+This implies that the time information operates independently from the primary network flow.
+Thus, we refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}$:
+$\mathcal{L}_{\text{TLA}}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||O(t;\mathbf{w}_{l})-\tilde{O}(t;\mathbf{w}_{l},\mathbf{s}_{l})||^{2}],$ (4)
+where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers.
+$O(t;\mathbf{w}_{l})$ is the intermediate activation of the full-precision model representing the ground truth, and $\tilde{O}(t;\mathbf{w}_{l},\mathbf{s}_{l})$ is the quantized activation.
+This objective function indicates that the chosen weight parameters are consistently updated across different time steps, so as to ensure robustness to diverse temporal inputs.
+Different from other methods [9, 15] that obtain different sets of quantization parameters for each time step, we only use a single set for varying time steps, improving time efficiency and memory storage.
+3.3.3 Critical Module Alignment
+While inaccurate time embedding quantization reduces performance under low-bit settings, it does not cause the complete generation failure observed in fully quantized models.
+Through careful layer-wise empirical study, we make the following observation:
+Property : Not all activations respond equally to reduced bit-width, as different activations exhibit varying levels of sensitivity, with certain critical layers being especially sensitive to quantization.
 Figure 3: Effect of decreasing different activations’ bit-width on the model performance.
 The generation failure of FeedForward layers emerges at 6 bits, while all other linear layers barely fail at 4 bits and all convolution layers only fail at 4 bits.
-Property : Different activations have different sensitivity towards decreasing bit-width.
-Low-bit activation quantization is often more difficult than weight quantization and remains the key challenge for diffusion model quantization.
-Compared with weights that can be categorized into linear layers and convolutional layers, activation types are more diverse and complicated.
-Figure 3 shows the sensitivity of different layers’ activations to quantization.
-Specifically, we quantize three different types of activations to lower bits while maintaining the others’ bit-width to 8-bit, and observe how the decreasing bit-width effects generation performance.
-Surprisingly, we observe that the FeedForward layer [8] activations cause generation failure at as large as 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits.
-This indicates that a small amount of activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.
-2.3 Quantization via Efficient Selective Finetuning
-We first theoretically discuss the potential reasons for the failure of previous methods and the underlying rationale for weight fine-tuning.
-Then we introduce a finetuning method QuEST for diffusion models that can significantly boost low-bit performance with less time and memory usage.
-Figure 1(c) shows our framework, being a distillation-based finetuning strategy consisting of selective weight optimization and network-wise scaling factor optimization.
-To further address the loss of temporal information due to quantization, an efficient time-aware activation quantizer is proposed.
-2.3.1 Sufficiency of Finetuning
-Motivated by the phenomenon that current PTQ methods completely fail when applied to low-bit diffusion models and driven by the phenomenon that the performance degradation caused by activation bit-width decrease is much more significant than weight bit-width (8-bit to 4-bit) in diffusion models, we would like to explore theoretically why previous PTQ methods fail and how to overcome this.
-We first review the underlying theory underpinning previous methods, which typically employ the reconstruction-based post-training quantization approach.
-Denote the diffusion model’s activations at time $t$ as $\mathbf{z}_{t}=[z_{1,t},z_{2,t},...,z_{n,t}]$, the task loss as $L(z_{n,t},\overline{z}_{n,t})$, where $n$ is the number of layers and $\overline{z}_{n,t}$ is the ground-truth.
+Fig.
+3 illustrates the sensitivity of different activations to quantization.
+Specifically, we quantize three different types of activations to lower bits while maintaining the others’ bit-width to 8-bit, and observe how the decreasing bit-width affects generation performance.
+Compared to weights that only fall into linear and convolutional layers, activations are more diverse and complex, making their effective quantization more challenging.
+Surprisingly, we observe that the FeedForward layer [7] activations cause generation failure at as early as 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits.
+This indicates that these activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.
+Denote $\mathbb{C}_{\text{A}}$ as the set containing all attention-related layers and given their image calibration inputs $z_{t,l}$, we optimize the corresponding weights and all quantization parameters $\mathbf{s}$, except for the ones already updated:
+$\mathcal{L}_{\text{CMA}}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||O(z_{t,l};\mathbf{w}_{l})-\tilde{O}(\tilde{z}_{t,l};\mathbf{w}_{l},\mathbf{\hat{s}})||^{2}],$ (5)
+where $\mathbf{w}_{l}$ are the weight parameters of the $l$th layer, $\tilde{z}_{t,l}$ is the quantized layer input, $\mathbf{\hat{s}}=\mathbf{s}\setminus\mathbf{s}_{l},l\in\mathbb{C}_{\text{TE}}$, which represents all the quantization parameters without the ones already finetuned in Eq.
+4.
+Note that we use different inputs to optimize each module, so as to enhance the robustness of the modules to the input perturbations.
+3.3.4 Progressive Alignment with Global Loss
+As investigated in the previous sections, two crucial types of layers are identified and selected for weight finetuning to enable quantization efficiency: time embedding layers and attention-related layers.
+We progressively align these components with the full-precision model due to their distinct, non-overlapping functionalities.
+Since temporal information is independent of the image input and determined early in the model, we first finetune the time embedding layers to provide accurate time step guidance for each subsequent module.
+Then we optimize the attention-related modules with the refined time embeddings.
+However, the above selective finetuning strategy only aligns the local information in the model, but is unaware of the global error reduction of the quantized model and the quantization parameters of the unselected layers.
+To improve the final generated images’ quality, we further aim to minimize the target task loss to provide global supervision:
+$\mathcal{L}_{\text{G}}=\mathbb{E}_{t}[||O(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s})||^{2}],$ (6)
+where $\mathbf{w}$ represents all the model weights, $O(x_{t};\mathbf{w})$ represents the final output of the full-precision counterpart and $\tilde{O}(x_{t};\mathbf{w},\mathbf{s})$ is the final output of the quantized model.
+By integrating Eq. 4, Eq. 5 and Eq. 6, the final objective is formulated as:
+$\mkern 28.0mu\arg\min_{\mathbf{w}_{l}}(\mathcal{L}_{\text{TLA}}+\mathcal{L}_{\text{CMA}}+2\mathcal{L}_{\text{G}}),\ \ l\in\mathbb{C}_{\text{TE}}\cup\mathbb{C}_{\text{A}}.$ (7)
+3.4 Finetuning from a Theoretical Perspective
+The above proposed method is motivated by the intuition that finetuning the model weights can adjust the activation distribution such that the imbalanced activation phenomenon can be alleviated.
+In this part, we attempt to explain why finetuning may be a feasible solution, offering additional insights for readers.
+However, we note that this is not a theoretical guarantee of the proposed method.
+We first review the underlying theory underpinning conventional post-training-quantization methods, which typically employ the reconstruction-based approach.
+Denote the full-precision diffusion model’s activations at time $t$ as $\mathbf{z}_{t}=[z_{1,t},z_{2,t},...,z_{n,t}]$, the final loss as $L(\mathbf{z}_{t};\mathbf{w})$, where $n$ is the number of layers.
 $L$ can be any loss function and here we use the mean squared error (MSE).
 We treat quantization as a type of perturbation and formulate the influence of activation quantization using Taylor expansion, assuming model weight $\mathbf{w}$ is frozen:
-$\mathbb{E}[L(z_{n,t}$ $+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]\approx\Delta^{T}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (3)
+$\mathbb{E}[L(z_{n,t}$ $+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]$
+$\approx\Delta^{\mathrm{T}}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{\mathrm{T}}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (8)
 where $\Delta$ is the activation perturbation, $\overline{\mathbf{g}}^{(\mathbf{z})}$ is the gradient and $\overline{\mathbf{H}}^{(z_{n,t})}$ is the Hessian matrix.
@@
 Thus the above equation can be simplified to:
-$\mkern-18.0mu\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta=\frac{1}{2}(\tilde{z}_{n,t}-z_{n,t})^{T}\overline{\mathbf{H}}^{(z_{n,t})}(\tilde{z}_{n,t}-z_{n,t}).\mkern-18.0mu$ (4)
-This leads to a rather straightforward conclusion: Quantization error can be reduced by reconstructing the model output.
-However, under low-bit settings, the reasoning from Equation 3 to Equation 4 is inaccurate, where the activation perturbation $\Delta$ is too large for meaningful Taylor expansion.
+$\mkern-18.0mu\frac{1}{2}\Delta^{\mathrm{T}}\overline{\mathbf{H}}^{(z_{n,t})}\Delta=\frac{1}{2}(\tilde{z}_{n,t}-z_{n,t})^{\mathrm{T}}\overline{\mathbf{H}}^{(z_{n,t})}(\tilde{z}_{n,t}-z_{n,t}).\mkern-18.0mu$ (9)
+However, under low-bit settings, the reasoning from Sec.
+3.4 to Eq.
+9 is inaccurate, where the activation perturbation $\Delta$ is too large for a meaningful Taylor expansion.
 Thus we have the following proposition:
@@
 Reconstruction-based post-training quantization methods may lose their theoretical guarantee due to the large value perturbations under low-bit quantization.
-To resolve the above problem, we can use the following theorem by transforming $\Delta$ into a smaller perturbation $\epsilon$:
+Since the inaccuracy arises from the large activation perturbation $\Delta$, we transform $\Delta$ into a smaller perturbation $\epsilon$ and derive the following theorem:
 Theorem 3.2.
-Given an $n$ layer diffusion model at time $t$ with quantized activations as $\tilde{\mathbf{z}}_{t}=[\tilde{z}_{1,t},\tilde{z}_{2,t},...,\tilde{z}_{n,t}]$ and $\tilde{z}_{n,t}=z_{n,t}+\Delta$, where $z_{n,t}$ is the ground-truth and $\Delta$ is the large perturbation caused by low-bit quantization.
-Denote the target task MSE loss as $L(\tilde{z}_{n,t},\overline{z}_{n,t})$, the quantization error can be transformed into:
+Given an $n$ layer diffusion model at time $t$ with quantized activations as $\tilde{\mathbf{z}}_{t}=[\tilde{z}_{1,t},\tilde{z}_{2,t},...,\tilde{z}_{n,t}]$ and $\tilde{z}_{n,t}=z_{n,t}+\Delta$, where $z_{n,t}$ is the ground truth and $\Delta$ is the large perturbation caused by low-bit quantization.
+Denote the target task MSE loss as $L(\mathbf{z}_{t};\mathbf{w})$, the quantization error can be transformed into:
 $\mathbb{E}[L(z_{n,t}+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]$
-$\approx$ $2\epsilon^{T}\sum_{i=1}^{K}(\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}-z_{\text{FP}})+\frac{1}{2}\sum_{i=1}^{K}(\tilde{z}_{n,t}^{i}-z_{\text{FP}})^{T}\overline{\mathbf{H}}^{(z_{n,t}+(i-1)\epsilon)}(\tilde{z}_{n,t}^{i}-z_{\text{FP}})$ (5)
-where $\mathbf{w}_{n}$ is the weight for layer $n$ and $\tilde{z}_{n,t}^{i}=\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}$, $\epsilon$ is a perturbation small enough for Taylor expansion, $K$ is a constant and $\Delta=K\epsilon$.
-Proof is provided in Appendix B.
-Theorem 2.2 indicates that: To minimize the quantization error, ideally we can finetune $\mathbf{w}_{n}$ so that for any $i$, the weights can fit to the data/activation $\tilde{z}_{n-1,t}^{i}+(i-1)\epsilon$ which the full-precision counterpart may not fit, making the model more robust to varying inputs.
-In other words, we are finetuning the model weights for better robustness towards large input activation perturbations, therefore enabling easier quantization.
-We also find that the second term in Equation 2.2 can be minimized simultaneously with the first term, so that the total quantization error can be reduced via a common goal: Activation output alignment via finetuning $\mathbf{w}_{n}$.
-2.3.2 Quantization via Selective Finetuning
-Motivated by the unraveled properties of quantized diffusion models and the shortcomings in existing quantization approaches, we propose an efficient finetuning strategy for low-bit diffusion model quantization.
-The components of our method are detailed below.
-Data-Free Partial-network-wise Training.
-To alleviate the need of substantial training data, we construct the calibration set in a data-free manner.
-By feeding sampled Gaussian noises $x_{T}$ into the full-precision model and sample iteratively over different time steps, we can obtain the calibration data needed for finetuning the quantized model.
-In practice, the number of samples generated for each time step is 128 or 256, indicating that we only have to inference the full-precision model a few times to obtain the needed number of calibration samples.
-As depicted in Figure 1(c), to overcome the quantization difficulty demonstrated in property efficiently, we update partial model weights ($\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$) that only account for a small subset of parameters related to the time step $t$.
-The other weight parameters (such as $\mathbf{W}_{l^{{}^{\prime}}}$, $\mathbf{W}_{l^{{}^{\prime\prime}}}$) are frozen during optimization.
-For example on LDM-4 [32], at most 7% of the parameters are included for adjustment.
-The choices for the weights to be finetuned are discussed in the following sections.
-To better conserve the performance and behavior of the full-precision model, we use both its intermediate activations and final outputs as the ground-truth for supervision.
-Figure 4 visualizes the generation results and demonstrates the effectiveness of this strategy in preserving image details.
-The activation quantization parameters can be viewed as additional model parameters, therefore we further propose a network-wise training strategy.
-Different from quantization methods using layer-wise or block-wise reconstruction [19, 34] that binds quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors for each time step together with the partial weight parameters.
-Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, we update the required parameters at once.
-In this way, we significantly save the time and memory needed for quantization.
-Note that the quantization parameters for the weights are not updated once initialized, in order to maintain memory efficiency during deployment.
-Efficient Time-aware Activation Quantization.
-As discussed in property , accurate time embedding quantization is critical.
-To discriminate between different time steps, we design a temporal quantizer specifically for activation quantization that adopts different sets of quantization parameters for different time steps.
-While these parameters consumes negligible memory and does not effect inference speed, learning independent parameters for each time step leads to immense time cost.
-Motivated by the fact that neighboring time steps yield similar functions in diffusion models [3], we cluster neighboring time steps so that they share the same quantization parameters and are calculated by the same calibration data $x_{t}$: $\mathbf{s}(t)=\mathbf{s}(t+1)=...=\mathbf{s}(t+\Delta t)$, where the time steps used for optimization are sampled uniformly, and each chosen time step’s statistics are used to represent $\Delta t$ adjacent ones.
-Though most time steps’ quantization parameters are approximated, this strategy saves the optimization time by $t$ times without compromising much performance.
-Section 3.2 further provides an analysis on the performance and efficiency trade-off.
-Table 3: Unconditional image generation results of LDM-4 models on LSUN-Bedrooms/Churches 256$\times$256. "N/A" denotes generation failure. "FP" denotes the full precision model.
+$\approx$ $2\epsilon^{\mathrm{T}}\sum_{i=1}^{K}(\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}-z_{\text{n,t}})$
+$+\frac{1}{2}\sum_{i=1}^{K}(\tilde{z}_{n,t}^{i}-z_{\text{n,t}})^{\mathrm{T}}\overline{\mathbf{H}}^{(z_{n,t}+(i-1)\epsilon)}(\tilde{z}_{n,t}^{i}-z_{\text{n,t}})$ (10)
+where $\mathbf{w}_{n}$ is the weight for layer $n$ and $\tilde{z}_{n,t}^{i}=\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}$, $K$ is a large constant and $\Delta=K\epsilon$.
 […]
-Selective & Progressive Layer Alignment.
-As investigated in Section 2.2, we identify two crucial types of activations for image generation performance and select them for weight finetuning: Time embeddings and attention-related activations (e.g.
-FeedForward layers in Stable Diffusion).
-We progressively update these components due to their distinct, non-overlapping functionalities.
-Figure 4: Generated images compared with Q-Diffusion under two bit-width types: W8A8 and W4A4, using LDM-4 for LSUN-Bedrooms.
-Our method not only succeed well in generating low-bit images, but also aligns more closely with the full-precision counterpart.
-In a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations.
-This implies that the time information operates independently from the primary network flow.
-Thus, we first refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}(t)$:
-$\mathcal{L}_{\text{TE}}^{t}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||\overline{O}(e(t);\mathbf{w}_{l})-\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}(t))||^{2}],$ (6)
-where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers, $e(t)$ is the fixed time embedding for time $t$.
-$\overline{O}(e(t);\mathbf{w}_{l})$ is the final activation output of the full-precision model representing the ground-truth, and $\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}(t))$ is the quantized output.
-This objective function indicates that the chosen weight parameters are consistently updated across different time steps, so as to ensure robustness to diverse temporal inputs.
-Nevertheless, only the small subset of scaling factors relevant to each specific time step are adjusted for temporal information discrimination.
-Apart from time embedding layers, attention-related layers are also vital (according to property ).
-As depicted in Figure 3, reducing the activation bit-width of these layers degrades model performance the most dramatically.
-Denote $\mathbb{C}_{\text{A}}$ as the set containing all these layers and given the image calibration samples $x_{t}$, we optimize the corresponding weights $\mathbf{w}_{l}$ and quantization parameters $\mathbf{s}(t)$.
-However, we exclude the quantization parameters already optimized in Equation 6:
-$\mkern-18.0mu\mathcal{L}_{\text{A}}^{t}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||\overline{O}(x_{t};\mathbf{w}_{:l})-\tilde{O}(x_{t};\mathbf{w}_{:l},\mathbf{s}(t)\setminus s_{\mathbb{C}_{\text{TE}}}^{t})||^{2}],\mkern-18.0mu$ (7)
-where $\mathbf{w}_{:l}$ are the weight parameters of the first $l$ layers. This strategy resembles linear probing, where only the layer directly associated with the activation output is finetuned.
-Furthermore, we aim not only for each independent layer’s quantization error reduction, but also for the final generated images’ quality improvement.
-Therefore, we adopt the Task Loss Guidance strategy where the target task loss $\sum_{t}\mathcal{L}_{d}^{t}$ is introduced and the ground-truth is approximated by the output of the full-precision counterpart.
-By integrating Equation 6 and Equation 7, the final objective is formulated as follows:
-$\mathcal{L}$ $=\sum_{t}((\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{d}^{t})+(\mathcal{L}_{\text{A}}^{t}+\mathcal{L}_{d}^{t}))$
-$=\sum_{t}(\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{\text{A}}^{t}+2\mathbb{E}[||\overline{O}(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s}(t))||^{2}]).$ (8)
+Table 3: Quantization performance on LSUN-Bedrooms/Churches 256$\times$256.
+“N/A” denotes generation failure.
+“*” denotes the results obtained by re-implementing the open-source code.
+More baseline and metric comparisons are included in the Appendix.
+Theorem 3.2 indicates that, to minimize quantization error, $\mathbf{w}_{n}$ should ideally be fine-tuned so that, for any $i$, the weights fit the corresponding input $\tilde{z}_{n-1,t}^{i}+(i-1)\epsilon$.
+This adjustment captures variations that the full-precision model may overlook.
+In other words, fine-tuning optimizes model weights for better robustness towards large input activation perturbations, facilitating easier quantization.
+Moreover, since the finetuned and quantized model is aligned with the original full-precision model, the potential impact on generation performance can be avoided.
+Note that the second term in Theorem 3.2 can be ignored within an acceptable upper bound, as it is of second order and shares a common zero-loss solution with the first term.
 4 Experiments
@@
 To verify the effectiveness of our proposed method, we conduct experiments on three types of generation tasks: Unconditional image generation on LSUN-Bedrooms and LSUN-Churches datasets [37], class-conditional image generation on ImageNet [4], and text-to-image generation.
-The model architectures we quantize include LDM and Stable Diffusion [32], and uses "WnAm" to represent the quantization setting: n-bit weight quantization and m-bit activation quantization.
+The model architectures we quantize include LDMs and Stable Diffusion [28], and use ”WnAm” to represent the quantization setting: n-bit weight quantization and m-bit activation quantization.
 DDIM samplers [13] are adopted for LDMs and the PLMS sampler [25] is used for Stable Diffusion.
-We generate 256 samples per time step for constructing the calibration set, and train 20 epochs over each time step.
-The Adam optimizer [18] is adopted and the learning rate for weight finetuning and scaling factor finetuning are set as $1e^{-5}$ and $1e^{-4}$ respectively.
-Table 4: Class-conditional generation results using LDM-4 on ImageNet 256$\times$256.
+We generate 256 samples per time step for constructing the calibration set.
+The Adam optimizer [17] is adopted and the learning rate for weight finetuning and scaling factor finetuning is set as $1e^{-5}$ and $1e^{-4}$ respectively.
 […]
+Table 4: Quantization performance on ImageNet 256$\times$256. “*” denotes the results obtained by re-running the open-source code.
 […]
@@
 Unconditional Generation: We evaluate the performance of our method over LDM-4 (LSUN-Bedrooms 256$\times$256) and LDM-8 (LSUN-Churches 256$\times$256) using the DDIM sampler with 200 and 500 time steps, respectively.
-Results are shown in table 3.
-The results are evaluated using FID, where the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet [19].
-Table 5: Text-to-image generation on Stable Diffusion (512$\times$512) using COCO2014 prompts.
+Results are shown in Tab.
+3 using FID, where our method outperforms the other baselines by a good margin.
+Note that the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet [18], thus not included.
+We further provide comparison with TFMQ-DM [15] in Appendix 6.
+[…]
+Table 5: Quantization performance on Stable Diffusion v1.4 (512$\times$512) using COCO2014 prompts.
 […]
 Class-conditional Generation: We evaluate the performance using LDM-4 on ImageNet 256$\times$256 using the DDIM sampler (20 steps).
-As shown in Table 4, three metrics are used for evaluation.
+As shown in Tab.
+4, three metrics are used for evaluation.
 Note that sFID uses additional intermediate spatial features for calculation compared with FID.
-We can also see that FID is not a valid metric for ImageNet LDM-4 evaluation: All methods have lower FID when quantized to lower-bits, conflicting with human perception.
-We show that our method not only succeeds in W4A4 quantization, but also significantly improve the generation quality on higher bit settings.
-Under W8A8, our method is able to achieve 2.82 improvement over sFID and 8.13 increase in Inception Score.
-Under W4A8, we improve the SOTA PTQ method by 1.57 sFID and 8.77 IS.
+We can also see that FID is not a valid metric for ImageNet LDM-4 evaluation: All methods have lower FID when quantized to lower bits, conflicting with human perception.
+We show that our method not only succeeds in W4A4 quantization, but also improves the generation quality under higher bit settings.
+Under all three kinds of bit-width settings, our method is able to outperform the SOTA PTQ methods and EfficientDM in both sFID and IS.
+Examples of our generated images are included in Appendix 11.
 Text-to-image Generation: We use Stable Diffusion v1.4 as the model for quantization with the PLMS sampler sampling 50 time steps.
-Table 5 shows the results.
+Tab.
+5 shows the results.
 Images are generated based on the 10,000 prompts sampled from the COCO2014 [23] validation set, and CLIP Score is calculated based on the ViT-B/16 backbone.
 Given the limited works done on Stable Diffusion, we can only compare with Q-Diffusion and the full-precision baseline.
-Table 6: Influence of task loss supervision on model performance.
-[…]
-Efficiency comparison and effects of components.
-Table 6 compares with and without the use of task loss $\mathcal{L}_{d}$.
-The results show that using the difference between the outputs of the quantized model and the full-precision counterpart for supervision is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for $\text{SLA}_{\text{TE}}$ and $\text{SLA}_{\text{A}}$, respectively.
-However, when the learning process is only supervised by the task loss, the performance degrades by 7.13 FID and 9.39 sFID for $\text{SLA}_{\text{TE}}$, indicating that the task loss alone is insufficient.
-Table 7 compares efficiency and performance between our method and the post-training quantization approach on the LSUN-Bedrooms dataset when using LDM-4.
-Our approach, while utilizing the same quantity of calibration data as the PTQ method, is more time-efficient and incurs only a 20% increase in GPU memory usage.
-We also show how each proposed component improves the generation performance.
-The results indicate that sequentially finetuning time embedding layers followed by finetuning attention-related layers leads to consistent performance improvements.
-Table 7: Component and efficiency comparisons on LDM-4 (LSUN-Bedrooms 256 $\times$ 256).
-Method
-Bit-width
-(W/A)
-Calibration
-data size
-Time cost
-(GPU hours)
-Memory cost
-Model size
-[…]
-FP 32/32 - - - 1045.6 2.95
-PTQ [19] 4/8 5120 23.08 10334 148.4 5.37
-Baseline 4/8 5120 11.52 9822 148.4 6.95
-+ $\text{SLA}_{\text{A}}$ 4/8 5120 13.13 12134 148.4 4.41
-+ $\text{SLA}_{\text{A}}$ + $\text{SLA}_{\text{TE}}$ 4/8 5120 15.25 12450 148.4 3.26
-Figure 5: Effect of uniformly sampled time steps number. Circle size represents quantization computation cost. Red square indicates the value we choose.
-Time step sampling frequency analysis.
-While time steps are sampled uniformly, will the performance consistently improve when we sample more frequently?
-Figure 5 illustrates the trade-off between the increased number of sampled time steps (’T-Num’) and generation performance.
-Metrics are calculated using 10,000 samples generated from LDM-4 on LSUN-Bedrooms.
-The results show that the though performance improves consistently, the amount of growth gradually decreases.
-Starting from 5 time steps, using 15 more time steps leads to 25.22 sFID improvement.
-But using another 30 time steps (from 20 to 50) only leads to 2.92 sFID improvement.
-Given the linearly increasing time cost associated with using more time steps, we opt for only 20 out of 200 time steps to enhance quantization efficiency.
-How QuEST influences model properties.
-Our approach is based on the certain characteristics observed in quantized diffusion models, hence we aim to analyze how our fine-tuning strategy counters these identified challenges.
-As shown in Table 2, finetuning time embedding layers leads to even better performance than the full-precision counterpart.
-We infer that this might be due to both the additional parameters introduced by quantization and the additional task loss supervision.
-Figure 2 further implies that our method adjusts the activation distribution so that it can be more easily quantized.
-We see that the ranges of the activation value distributions shrink from [-10, 34] to [-4, 14] and [-11, 20] to [-4, 4].
-The standard deviations also decrease from 0.171 to 0.157 and from 0.073 to 0.071.
-This indicates a higher concentration of values around the mean, which can effectively reduce both rounding and clipping errors.
+Figure 4: Visual comparison with Q-Diffusion and EfficientDM. QuEST outperforms the baselines with better visual quality.
+4.3 Ablations and Discussions
+Efficiency comparison with PTQ methods and the impact of individual components.
+Tab.
+7 compares the efficiency and performance against the post-training quantization (PTQ) approach on the LSUN-Bedrooms dataset.
+Although our method uses the same amount of calibration data as the PTQ approach, it achieves better time efficiency with only a 20% increase in GPU memory usage.
+We also illustrate the contribution of each component to generation performance.
+The results indicate that sequentially finetuning the time embedding layers, followed by attention-related layers, yields consistent performance improvements.
+Tab.
+7 presents a comparison of performance with and without the global loss $\mathcal{L}_{\text{G}}$.
+The results indicate that supervising the quantized model using the output difference from the full-precision counterpart is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for TLA and CMA, respectively.
+However, when the learning process is only supervised by the global loss, we find that the performance degrades by 7.13 FID and 9.39 sFID for TLA, suggesting that the global loss alone is insufficient for optimal performance.
+Bit-width Method Time (h) Memory (MB) Iters FID $\downarrow$ EfficientDM 2.60 12004 32k 15.15 Full-finetune 0.85 15076 2.2k 5.38 W4A8 Ours 0.45 12178 2.2k 3.82 EfficientDM 2.60 12004 32k 10.60 Full-finetune 0.85 15076 2.2k 6.36 W4A4 Ours 0.45 12178 2.2k 6.12
+Table 8: Efficiency comparison with other finetuning methods.
+How QuEST adjusts the activation distribution.
+Our approach is motivated by the imbalanced activation distribution in diffusion models, hence we aim to analyze how our fine-tuning strategy addresses this challenge.
+As shown in Fig.
+2, our method refines the activation distribution, making it more conducive to quantization.
+Specifically, the activation value ranges shrink from [-10, 34] to [-4, 14] and from [-11, 20] to [-4, 4].
+Additionally, the standard deviations decrease from 0.171 to 0.157 and from 0.073 to 0.071, while the mean remains consistent.
+This results in a more compact activation distribution, effectively reducing both rounding and clipping errors during quantization.
+Comparison with precomputed time embeddings.
+In diffusion models, time embeddings are independent of input conditions and noise.
+A potential approach is to precompute these embeddings and reuse them directly.
+However, this strategy overlooks the compatibility between different modules in a quantized model.
+We take this into consideration and optimize the time embeddings with $\arg\min_{\mathbf{w}_{l}}(\mathcal{L}_{\text{TLA}}+\mathcal{L}_{\text{G}}),\ \ l\in\mathbb{C}_{\text{TE}}$ so that the time embedding layers are also trained to minimize the final prediction error.
+As shown in Tab.
+2, adding this optimization objective enhances quantization performance, even surpassing the full-precision baseline (which uses precomputed features).
 Integration with LoRA finetuning.
-Finetuning quantized models is not trivial.
-Thus, we use QALoRA [11] that extends LoRA [16] and finetune on the ImageNet 256$\times$256 dataset.
-The weight rank of 32 is used for the LoRA modules and the parameters are trained over 100 time steps for 160 epochs.
-We find that integrating the QALoRA technique leads to a 5.62 decrease in FID, indicating that finetuning the original layers is a better solution for performance preservation.
+Different ways exist for finetuning quantized models.
+We further employ QALoRA [9] to finetune on the ImageNet 256$\times$256 dataset.
+A rank of 32 is used for the LoRA weights, and the parameters are trained over 100 time steps for 160 epochs.
+We find that integrating the QALoRA technique leads to a 5.62 increase in FID, indicating that finetuning the original layers is a better solution for performance preservation.
+Efficiency comparison with other finetuning methods.
+We compare with EfficientDM and full-finetuning in terms of actual training costs on LDM-4 in Tab.
+8.
+The setting of full-finetuning is aligned with our method.
+We observe that: compared with EfficientDM, our method requires fewer training iterations and time to obtain better performance with comparable GPU memory cost.
+Compared with full-finetuning, our method costs less time and memory, as well as achieving better performance.
+The bottleneck in computational costs becomes more severe when scaled to larger models such as Stable Diffusion.
+We find that while full-finetuning quickly encounters OOM, our method is able to finetune SD on a single GPU with 48GB memory.
 5 Conclusion
 We have proposed QuEST, an efficient data-free finetuning framework for low-bit diffusion model quantization.
-Our method is motivated and guided by the three underlying properties found in quantized diffusion models.
-We also theoretically demonstrate the sufficiency of finetuning, interpreting it as a way to enhance model robustness against large activation perturbations.
+Our method is motivated by the current challenge in low-bit diffusion model quantization and guided by the two underlying properties found in quantized diffusion models.
 To alleviate the performance degradation, we propose to finetune the time embedding layers and the attention-related layers under the supervision of the full-precision counterpart.
 Experimental results on three high-resolution image generation tasks (including Stable Diffusion) demonstrate the effectiveness and efficiency of QuEST, achieving low-bit compatibility with less time and memory cost.
-Appendix A Related Works
-A.1 Diffusion Model Inference
-Diffusion models [15, 32] generate samples via a learned iterative denoising process.
-During inference, the initial input is sampled from a Gaussian distribution: $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the final output $x_{0}$ is obtained by removing noises from $x_{T}$ subsequently using:
-$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\tilde{\boldsymbol{\mu}}_{\theta,t}(x_{t}),\tilde{\beta_{t}}\mathbf{I}).$ (9)
-In this Markov Chain, $\tilde{\boldsymbol{\mu}}_{\theta,t}$ and $\tilde{\beta_{t}}$ are calculated from the model’s output.
-For a typical diffusion model, the denoising stage requires hundreds to thousands of iterations.
-This laborious sequential process makes efficient inference and high-resolution generation extremely challenging.
-Latent Diffusion Models (LDMs) [32] were developed to overcome the challenges.
-They encode the data into a latent space and then apply the diffusion process within this space, enabling high-resolution image generation with fewer time steps.
-The output images are produced by decoding the latent representations.
-LDM’s concept and network quantization complement each other, mutually enhancing the efficiency of diffusion models.
-Practically, diffusion models typically adopt a UNet architecture [33], with some modules substituted by attention blocks.
-In a latent diffusion model, an encoder and a decoder are incorporated for projecting images into latent representations and subsequently generating images from the latent space, respectively.
-Usually encoders and decoders are lightweight and computationally inexpensive, so our focus is on quantizing the UNets in LDMs, in alignment with other works (as described below).
-A.2 Diffusion Model Quantization
-Model quantization is a dominant technique for optimizing the inference memory and speed of deep learning models by reducing the precision of the tensors used in computation.
-Few researches were done for quantizing diffusion models and they fall into two mainstream frameworks: Quantization-Aware Training (QAT) [17, 20, 21, 38] and Post-Training Quantization (PTQ) [22, 25, 28, 36, 37].
-QAT methods such as Q-DM [20] consider value precision as an additional optimization constraint, training all parameters from scratch with the target task objective.
-While it is effective for low-bit quantization, it is even more resource-intensive than training a full-precision model.
-Efficient-DM [11] further uses Low-rank adapters (LoRA) [16] to reduce training cost.
-However, it introduces extra weight parameters and still requires substantial training iterations.
-Post-Training Quantization methods aim to calibrate the quantization parameters based on a small calibration set.
-Pioneering works such as PTQ4DM [34] and Q-Diffusion [19] focus on sampling the outputs of the full-precision model under different time steps with certain probabilities to construct the calibration set.
-ADP-DM [35] partition time steps into different groups and optimize each group with differentiable search algorithms.
-PTQ-D [12] decomposes the quantization noise and correct them individually with full-precision model statistics.
-However, these methods generally fail under 4-bit or lower bit-width.
-To enable low-bit compatibility with high efficiency, our method adopts the general framework of PTQ but further introduces an efficient weight fine-tuning strategy.
-Appendix B Proof for Theorem 2.2
+Supplementary Material
+The supplementary material is organized as follows: Sec.
+6 provides comparison with TFMQ-DM; Sec.
+7 provides comparison on the low-resolution dataset; Sec.
+8 provides the proof and detailed analysis for Theorem 3.2; Sec.
+9 presents additional examples of the imbalanced distributions across different models; Sec.
+10 highlights the importance of the large values in activations; Sec.
+11 offers further generated examples from our method across varying bit-widths; and Sec.
+12 discusses limitations and broader considerations.
+6 More Baseline Comparisons
+We further compare with TFMQ [15] below:
+Bedroom W8A8 W4A8 TFMQ-DM 3.14 3.68 QuEST 3.03 3.26 ImageNet W8A8 W4A8 TFMQ-DM 10.79 10.29 QuEST 10.43 8.48
+Table 9: Comparing TFMQ.
+We also supplement the metrics for Table 3:
+W8A8 sFID $\downarrow$ IS $\uparrow$ QDiffusion 8.19 2.25 PTQD 9.89 2.25 EfficientDM N/A N/A Ours 6.86 2.27 W4A4 sFID $\downarrow$ IS $\uparrow$ QDiffusion N/A N/A PTQD N/A N/A EfficientDM 15.15 2.27 Ours 7.82 2.26
+Table 10: Additional metrics on LSUN-Bedrooms. “N/A” represents generation failure.
+7 Low-resolution dataset comparison
+We further include experiments on CIFAR10 in Tab. 11.
+W8A8 W4A4 Q-Diffusion 3.75 N/A EfficientDM 3.75 10.48 QuEST 3.71 9.37
+Table 11: FID comparison on CIFAR10.
+8 Proof for Theorem 3.2
 We provide the detailed proof for Theorem 3.2 here. The notations are consistent with the ones in the main paper.
 Since the perturbation $\Delta$ is too large for accurate Taylor expansion, we can resolve it by introducing a new perturbation $\epsilon=\Delta/K$, where we divide $\Delta$ by a constant $K$ so that $\epsilon$ is small enough for approximation.
-Then, Equation 3 is rewritten as follows:
+Then, Sec.
+3.4 is rewritten as follows:
 $\mathbb{E}[L(z_{n,t}+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]$
@@
 where the approximation step follows Taylor expansion and only the first two main components are kept.
-The first term in Equation B cannot be ignored because samples such as $z_{n,t}+(i-1)\epsilon$ may not be included in the learned distribution of the model.
+The first term in Sec.
+8 cannot be ignored because samples such as $z_{n,t}+(i-1)\epsilon$ may not be included in the learned distribution of the model.
 The second term can still be minimized by reconstruction since only the difference between quantized model output and ground-truth matters.
@@
 Figure 5: Illustrations of imbalanced activation distributions on conditional LDM4 (ImageNet 256$\times$256) and unconditional LDM4 (LSUN-Bedrooms 256$\times$256).
+Figure 6: Comparison of different corruptions made on different tokens.
 Given the objective function (MSE loss) of diffusion models, we analyze that:
@@
 Ground-truth $\overline{z}_{n,t}$ can be approximated by the full-precision output $z_{\text{FP}}$.
-We see that $\tilde{z}_{n-1,t}^{i}$ and $z_{\text{FP}}$ cannot be changed, thus to minimize Equation B, we need to finetune $\textbf{w}_{n}$.
-From a general perspective, Equation B also indicates that the model has not converged well to a local minimum given the perturbed inputs, thus when we finetune the model layers given the quantized inputs, we are actually training the model towards convergence over new samples and increasing its robustness.
-Appendix C Examples of Imbalanced Activation Distributions
-Apart from Figure 2, we show that the imbalance in the activation distribution is a common phenomenon in different model structures and datasets.
-In Figure 6, we show more results of activation distributions of latent diffusion models on ImageNet 256 $\times$ 256 and LSUN-Bedrooms 256 $\times$ 256.
-Appendix D Importance of large values in activations
-As shown in Figure 2, quite a few values are rather large and diversely distributed.
+We see that $\tilde{z}_{n-1,t}^{i}$ and $z_{\text{FP}}$ cannot be changed, thus to minimize Sec.
+8, we need to finetune $\textbf{w}_{n}$.
+From a general perspective, Sec.
+8 also indicates that the model has not converged well to a local minimum given the perturbed inputs, thus when we finetune the model layers given the quantized inputs, we are actually training the model towards convergence over new samples and increasing its robustness.
+9 Examples of Imbalanced Activation Distributions
+Apart from Fig.
+2, we show that the imbalance in the activation distribution is a common phenomenon in different model structures and datasets.
+In Fig.
+5, we show more results of activation distributions of latent diffusion models on ImageNet 256 $\times$ 256 and LSUN-Bedrooms 256 $\times$ 256.
+10 Importance of large values in activations
+As shown in Fig.
+2, quite a few values are rather large and diversely distributed.
 These values pose difficulties on activation quantization, and being rather important and not negligible.
@@
 The corruption is done by setting the token values as all zeros.
-As shown in Figure 7, we compare two settings: (1) corrupt a certain number of tokens randomly; (2) corrupt the same number of the tokens with the largest values.
-We see that when corrupting randomly, generation performance is hardly effected.
-However, corrupting the same amount of tokens (even only one token) with the largest values leads to significantly degenerated images.
-Figure 7: Comparison of different corruptions made on different tokens.
-Appendix E Generated image examples for different tasks
-E.1 Unconditional Image Generation
-The generated images for LDM-4 on LSUN-Bedrooms 256$\times$256 under different bit-widths are shown in Figure 8. Images for LSUN-Churches 256$\times$256 are shown in Figure 9.
+As shown in Fig.
+6, we compare two settings: (1) corrupt a certain number of tokens randomly; (2) corrupt the same number of the tokens with the largest values.
 (a) Full Precision
@@
 Figure 7: Unconditional image generation examples for LSUN-Bedrooms 256$\times$256.
+We see that when corrupting randomly, generation performance is hardly effected.
+However, corrupting the same amount of tokens (even only one token) with the largest values leads to significantly degenerated images.
+Figure 8: Text-to-image generation results on Stable Diffusion.
+11 More generated image examples
+11.1 Unconditional Image Generation
+The generated images for LSUN-Bedrooms 256$\times$256 under different bit-widths are shown in Fig. 7. Images for LSUN-Churches 256$\times$256 are shown in Fig. 9.
 (a) Full Precision
@@
 Figure 9: Unconditional image generation examples for LSUN-Churches 256$\times$256.
-E.2 Class-conditional image generation
-Figure 10 shows the generated images for 3 different classes.
+11.2 Class-conditional image generation
+Fig. 10 shows the generated images for 3 different classes.
 (a) Full Precision
@@
 Figure 10: Conditional image generation results for ImageNet 256$\times$256.
-Figure 11: Text-to-image generation results on Stable Diffusion.
-E.3 Text-to-image generation
-Figure 11 shows the generated images for two different prompts using Stable Diffusion v1.4 under different bit-width.
-Appendix F Limitations and Broader Impacts
+11.3 Text-to-image generation
+Fig. 8 shows the generated images using Stable Diffusion v1.4 under different bit-width.
+12 Limitations and Broader Impacts
 The primary objective of this paper is to further the research in enhancing the efficiency of diffusion models.